A Complete Bibliography of Publications in the SIAM Journal on Applied Mathematics: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 1.09

Title word cross-reference

2 [317, 332]. 3 [365]. K [58]. keff [164]. M [48]. R^{n} [410].

-Cones [317]. -Core [58]. -Polygons [48].

19 [340].

Absorbing [260, 237]. Absorption [398, 265, 37]. Abstract [3, 133]. Acceleration [128, 165]. Accounting [128]. Accumulation [365, 381]. Accurate [119]. Acoustic [356, 234, 187, 114, 260, 346, 237, 374, 276, 103]. Acoustical [401]. Acoustically [347]. Acoustically-Modulated [347]. Acoustics [77, 298]. Across [163]. Action [87]. Active [309, 182]. Adaptation [109]. Adapting [258]. Adaptive [30, 55]. Adder [60]. Adding [79, 62]. Additive [324]. Adhesion [20]. Aedes [44]. Aegypti [44]. Affect [98]. After [251]. Age [17, 375]. Age-Structured [17, 375]. Agents [148]. Aggregation [177]. Aging [85]. Agreement [291]. Agro [378]. Agro-Ecological [378]. Albopictus [44]. Algorithm [119, 400]. Alignment [291]. Allee [139, 82]. Allocation [171]. Alloys [112]. Almost [315]. Along [215]. Ambient [156]. Ampère [345]. Amplification [95]. Amplitude [111]. Anaerobic [178]. Analysis [297, 95, 42, 238, 186, 410, 348, 46, 26, 8, 138, 197, 359, 56, 323, 13, 77, 253, 183, 174, 235, 150, 185, 37, 153, 39, 178, 411, 414, 35, 89, 272, 182, 81, 408, 152]. Analytical [394]. Anchoring [104]. Angled [303]. Animal [117]. Anisotropic [190, 110, 65]. Anisotropy [349, 113]. Apodizer [376]. Application [255, 389, 337]. Applications [267, 317, 44, 264, 270, 81]. Applied [36]. Approach [396, 162, 53, 64, 387, 215, 303, 275, 225, 212]. Approximate [287, 236, 1, 91, 103]. Approximated [282]. Approximating [9]. Approximation [331, 348, 216, 332]. Arbitrary [5]. Architecture [49]. Arising [42, 201]. Array [77, 94]. Arrival [337]. Artificial [300, 240]. Aspects [263]. Assessment [260]. Asymmetric [352]. Asymptomatic [277]. Asymptotic [52, 213, 57, 201, 138, 220, 359, 50, 134, 336, 272, 154, 366, 90]. Asymptotically [171]. Asymptotics [6, 16, 102, 351]. Atomistic [262]. Attached [355]. Attenuation [142]. Attractors [106]. Autocatalytic [194]. Autonomous [367]. Averaging [255]. Avoiding [214]. Aw [83]. Aw-Rascle [83]. Axisymmetric [355]. Axon [403].

Bacterial [176]. Baking [197]. Balance [175]. Balanced [12]. Bank [407]. Basal [27]. Based [162, 161, 64, 13, 390, 290, 212, 107]. Basic [57, 90, 404]. Basins [141]. Basis [412, 312]. Batteries [266]. Bayesian [373]. BCC [406]. Beavers [282]. Becker [289]. Bed [172, 97, 366]. Bee [40]. Before [251]. Behavior [53, 411]. Behaviors [360]. Benthic [352]. Between [354, 352, 194, 143]. Beyond [385, 52]. Biaxial [390]. Bifurcation [62, 316, 81, 408]. Bifurcations [269, 363]. Biharmonic [231]. Bikerman [92]. Bimodal [266, 62]. Binding [308]. Binomial [61]. Biochemical [372]. Biofilm [78]. Biological [116, 177]. Bioproducing [335]. Biphasic [95]. Blade [235]. Blebbing [285]. Blocking [97]. Blood [328]. Blowup [111]. Bodies [182]. Body [215, 157]. Boltzmann [395]. Bonded [33]. Bone [278]. Boreholes [15]. Borne [320]. Bose [254]. Bound [384, 298, 41]. Bound-Preserving [384]. Boundaries [292, 107]. Boundary [210, 282, 331, 93, 301, 73, 174, 299, 296, 168, 260, 237, 51, 196, 355]. Bounded [20, 128]. Boundedness [205]. Bounds [100, 5]. Bragg [376]. Brain [169]. Branch [251]. Break [6]. Breaking [322]. Breakup [126]. Bridges [305]. Brine [302]. Brinkman [297]. Brownian [309]. Bubble [16, 4]. Bubbles [245]. Building [79]. Bulk [69]. Bump [14]. Burgers [232].

C [160]. C-Looping [160]. Cable [144]. Cahn [104, 99, 272]. Canard [280]. Capillary [141]. Capture [161]. Car [310, 327]. Car-Following [310, 327]. Carbon [382]. Carlo [386]. Case [71, 298]. Cavitating [275]. Cavitation [9]. Cavity [9]. Cell [43, 27, 61, 154, 285, 7, 328, 408, 246]. Cells [207, 32]. Cellular [60]. Centrality [239, 342]. Certain [101]. certainty [281]. Chains [389]. Change [117]. Changing [40, 109]. Channel [269]. Channels [134]. Characterization [12, 190]. Characterizations [326]. Charge [45, 86, 345, 151]. Charge-Group [151]. Chemical [79, 360, 389, 78]. Chemistry [130]. Chemostat [274]. Chemotactic [40]. Chemotherapy [169]. Chirality [209]. Circular [179]. Class [42, 219, 185, 205]. Classical [239, 342, 86]. Classifying [304, 113]. Climate [117]. Closure [216]. Cloud [34]. Co [122]. Co-Flow [122]. Coarse [273]. Coarse-Grained [273]. Coated [121, 65]. Coefficient [40, 248, 208, 124]. Coefficients [86]. Collapsing [377]. Collision [327]. Collisions [214]. Colloidal [23]. Colonies [40, 192]. Column [37]. Comb [38]. Combustion [47]. Common [350].

Communication [156]. Community [64, 173, 58]. Compartmental [43]. Competition [230, 193, 354, 44, 204, 352]. Competition-Diffusion [193]. Competitors [413, 44]. Complex [188, 12, 232, 394, 241, 89]. Complex-Balanced [12]. Complex-Frequency [89]. Component [46]. Composites [370]. Compressible [339]. Computation [100, 105]. Computational [372]. Computed [88]. Condensates [254]. Condensation [198]. Condition [282, 385, 268, 196]. Conditional [120]. Conditions [104, 73, 260]. Conductivity [370, 65]. Conductors [73]. Cones [317]. Confinement [316]. Conformal [299]. Congestion [405]. Conjugacies [363]. Connected [249, 37, 369, 298]. Conservation [210]. Consistency [25]. Consortium [335]. Constant [65]. Constitutive [353]. Constraint [27]. Construction [260]. Contact [21, 258, 149]. Contagion [323, 185]. Continuity [384]. Continuous [328]. Continuum [298, 212, 51, 41]. Contour [221]. Contrast [356, 220, 303, 412, 247]. Control [66, 314, 43, 340, 378, 382, 129, 335, 127, 149, 319, 81]. Control-Theoretic [66]. Controlling [129, 320]. Convergence [201, 317, 9]. Convergent [134]. Convexity [166]. Convolutional [115]. Core [65, 58]. Coronary [288]. Correlations [330, 195]. Cortex [49, 285]. Costs [295]. Coulomb [195]. Countercurrent [366]. Couple [376]. Coupled [374, 102, 69]. Coupling [338, 285, 107]. COVID [340]. COVID-19 [340]. Crime [66]. Criterion [412]. Critical [71]. Criticality [198]. Cross [11]. Cross-Diffusion [11]. Crowd [53]. Cryopreservation [32]. Crystal [242]. Crystals [93, 135, 151, 316]. Culling [149]. Culture [37]. Current [141, 67, 45]. Curvature [19]. Curves [263, 234].

Cyanobacteria [56]. Cyclic [199]. Cyclicity [413]. Cytoneme [161]. Cytoneme-Based [161].

D [365, 332]. Dam [6]. Dam-Break [6]. Damped [202]. Darwin [191]. Data [61, 358, 412, 103]. Data-Driven [412]. Deal [71]. Debt [407]. Debye [92]. Decaying [28]. Decentralized [306]. Decomposition [213, 131]. Deep [146, 97]. Deep-Frying [146]. Defect [242]. Defects [261]. Deffuant [348]. Deficiency [400]. Deformation [7, 217]. Degenerate [62]. Delay [199, 201, 310, 227, 408]. Delayed [68, 340, 319]. Delivery [85]. Delocalization [254]. Dendritic [43, 273, 408]. Dense [223, 4]. Density [27, 371]. Dependent [229, 5, 394, 165]. Depletion [331, 289]. Depth [5]. Depth-Dependent [5]. Derivation [283, 414, 246]. Derivatives [175]. Described [191]. Describing [308]. Description [246]. Design [376, 281]. Design-Structured [281]. Destabilization [106]. Detail [144]. Detailed [12]. Detailed-Balanced [12]. Detecting [222]. Detection [64]. Deterioration [362]. Determinacy [82]. Determining [180]. Deterministic [387]. Developing [32]. Development [255, 251]. Deviations [324]. Dewetting [75]. Diapause [68]. Dielectric [209, 21]. Difference [36]. Different [354]. Differentiability [3, 133]. Differentiable [363]. Differential [22, 199, 201, 34, 290]. Differentiation [180]. Diffraction [123, 303, 276, 341]. Diffuse [285]. Diffusing [102]. Diffusion [349, 68, 125, 365, 392, 11, 314, 193, 333, 26, 236, 313, 283, 136, 265, 86, 364, 102, 274, 69, 411, 179, 200, 130, 268, 311, 205, 182, 169, 19]. Diffusions [324]. Diffusive [120, 127, 307, 227]. Digestion [178]. Dimension [5]. Dimensional [67, 269, 370, 356, 271, 38, 138, 76, 219, 358, 69, 62, 89, 200, 268, 196, 351]. Dimensions [166, 144, 75, 159]. Dirac [393]. Direct [216]. Directed [139]. Directional [180, 161]. Disc [33]. Disconnection [51]. Discourse [211]. Discrete [358, 218, 117, 205]. Discrete-Time [218, 117]. Discretization [219]. Disease [173, 307, 81]. Diseases [277, 337, 320]. Disk [179]. Dislocation [107]. Dislocations [112]. Disorder [254]. Dispersal [155, 98]. Dispersed [370]. Dispersion [186]. Displacement [54]. Dissipation [367]. Distinguishability [5]. Distributed [199]. Distribution [183, 198]. Distributions [394, 266]. DNA [95]. Does [98]. Dollar [407]. Domain [162, 365, 383]. Domains [2, 181, 20, 357, 369]. Döring [289]. Double [303]. Doubly [291]. Dough [146]. Down [8]. Drawing [359]. Drift [200]. Drift-Diffusion [200]. Drive [38]. Driven [232, 158, 8, 412, 273, 117, 19]. Drop [217]. Droplet [126]. Drug [189, 85]. Drug-Eluting [189]. Dry [387]. Duct [237]. Due [47]. During [344, 362]. Dyadic [312]. Dynamical [400, 215]. Dynamics [396, 53, 255, 125, 407, 377, 267, 95, 348, 192, 138, 211, 56, 372, 194, 334, 59, 55, 256, 327, 149, 71, 274, 165, 69, 345, 307, 227, 319, 51, 285]. Dynamo [100].

Ecological [230, 378]. Eddy [67]. Edges [358]. Effect [139, 116, 85, 82, 274, 126]. Effected [351]. Effective [17, 270]. Effects [349, 277, 120, 59, 374, 84, 244]. Efficiency [349]. Efficient [119, 12, 278, 306]. Efficiently [376]. Eigenfunctions [240]. Eigenvalue [42, 73, 89, 351]. Eigenvalues [100, 228, 264]. Einstein [254]. Elastic [33, 190, 245, 409, 383, 50, 339, 7, 355]. Elasticity [228, 270]. Elastodynamics [224]. Elastography [1, 91]. Electric [135, 361]. Electric-Field-Induced [135]. Electrical [258, 5, 25]. Electro [21]. Electro-Wetting [21]. Electroanalytical [130]. Electrode [258, 197, 266]. Electrodiffusive [163]. Electrohydrodynamics [217]. Electromagnetic [209, 240, 45, 121, 347]. Electrons [191]. Electrostatic [38]. Electrostatics [151]. Element [235]. Elementary [243, 363]. Elimination [43]. Eluting [189]. Embedded [185, 270]. Embryonic [160]. Emergence [140]. Emission [382, 124]. Endemic [218]. Energy [104, 304, 28, 295, 275, 96]. Engineering [172]. Enhanced [163]. Enhancing [284]. Entropy [112, 281]. Envelope [381]. Environment [57, 90]. Environmental [66]. Environments [230, 106, 109, 246]. Epicardial [288]. Epidemic [344, 314, 57, 90, 206, 411, 311]. Epidemics [129, 325]. Epidemiology [201]. Epidermis [27]. Equation [257, 119, 144, 22, 199, 224, 164, 232, 108, 170, 271, 248, 379, 208, 124, 191, 92, 142, 299, 357, 364, 272, 268, 48]. Equations [297, 229, 384, 392, 199, 201, 24, 34, 301, 13, 313, 315, 399, 283, 187, 3, 133, 191, 216, 286, 395, 134, 290, 353, 36, 312]. Equilibria [76]. Equilibrium [10]. Equivalent [331]. Erratum [342, 90, 181, 91, 133]. Error [259]. Escape [43]. Essentially [267]. Estimates [259]. Estimating [281]. Estimation [137]. Euler [111]. Eulerian [339]. Evaporation [381]. Every [315]. Evolution [125, 186, 3, 133, 404, 48]. Evolving [23]. Exact [6, 296, 97, 134]. Exchange [407]. Excitation [298, 41]. Exclusion [14]. Exercise [23]. Existence [125, 105, 138, 290]. Expansions [52, 220]. Experiment [383]. Explosions [280]. Exponential [16, 298, 184]. Expression [394, 322]. Extended [45, 380]. Extinction [218]. Eye [325].

Factorization [346, 103]. Failure [241]. Family [263, 184]. Fangcang [340]. Fano [94]. Far [121, 286, 409]. Far-Field [121, 286]. Fast [255, 191, 120]. Federated [306]. Fermentation [290]. Fermi [332]. Ferroelectric [72]. Ferronematics [269]. Fiber [376, 359]. Fibrin [251]. Fibrous [246]. Field [53, 100, 72, 348, 116, 135, 388, 323, 121, 129, 28, 286, 409, 65, 250, 69, 212, 35, 103]. Fields [226]. Filament [145, 48]. Filippov [319]. Filled [353]. Film [8, 353]. Films [72]. Filter [147]. Filtration [279, 97]. Finding [400]. Fingering [33]. Fingers [343]. Finite [384, 97]. Finite-Volume [384]. Fins [296]. First [73, 302, 120]. First-Year [302]. Fishing [294]. Fixed [172, 341, 182]. Fixed-Bed [172]. Flight [410]. Flocks [214]. Floquet [396]. Flow [349, 167, 232, 350, 8, 122, 132, 249, 111, 128, 28, 397, 237, 252, 126, 414, 366, 4, 7]. Flows [213, 317, 338, 371]. Fluctuations [170, 371]. Fluid [210, 233, 253, 110, 28, 381]. Fluid-Infused [381]. Fluid-Particle [28]. Fluidic [339]. Fluids [297]. Flux [216]. Follicles [255]. Following [310, 327]. Foot [307]. Foot-and-Mouth [307]. Foraging [410]. Forest [386]. Form [413]. Formal [186]. Formation [214, 158, 93, 47, 402]. Formations [314]. Formed [113]. Forms [363]. Formulas [369]. Formulation [45, 397, 107, 312]. Formulations [331, 72]. Foundation [61]. Fourier [333]. Fractional [283, 149, 272]. Fractional-Order [149]. Fracture [262]. Fractures [96]. Fragmentation [127]. Frame [390]. Framework [259, 262, 32, 372, 345, 375]. Free [210, 155]. Free-Boundary [210]. Freezing [32]. Frequency [383, 346, 89, 341]. Frequency-Domain [383]. Friction [387, 351]. Friction-Type [351]. Front [186, 122, 337]. Fronts [8, 102]. Frying [146]. Full [162, 119]. Fully [351]. Functional [104, 49, 280]. Functionals [324]. Functions [207, 287, 358, 312].

Gap [17, 219]. Gardner [224]. Gas [374, 366]. Gas-Solid [366]. Gathering [148]. Gel [377]. Gelation [251]. Gene [394, 322]. General [155, 63, 329, 65, 395, 39, 336, 109]. Generalized [83, 170, 326, 92, 369, 58, 89, 196, 81]. Generation [189]. Generational [164]. Generations [41]. Generic [317]. Geological [141]. Geometric [33]. Geometrical [240]. Geometry [108, 134]. Geothermal [15]. Gierer [138]. Glioblastoma [11]. Global [30, 280, 319]. Glucagon [238]. Glucose [238]. Goldbeter [52]. Gradient [216, 364, 338]. Grain [292, 51, 107]. Grained [273]. Graph [386, 350, 313, 174, 184]. Graphene [361]. Graphical [279]. Graphons [321, 223]. Graphs [223]. Grating [376, 341]. Gravity [141]. Green [119, 312]. Grid [207]. Ground [254]. Group [151]. Groups [326]. Growth [68, 11, 137, 47, 168, 339, 169]. Guided [353].

Habitat [193, 117]. Hairs [183]. Half [293, 298]. Half-Plane [298]. Half-Space [293]. Hamiltonian [350]. Handling [80]. Haptotaxis [74]. Hard [195]. Hard-sphere [195]. Harmonic [15, 121, 234, 326, 260, 41]. Having [65]. Haystack [287]. Heart [160]. Hele [16]. Helioseismology [119]. Helmholtz [301, 77, 142, 357]. Heterogeneity [155, 266, 362, 140]. Heterogeneous [57, 90, 136, 260]. Hexagonal [292]. Hidden [50]. Hierarchical [59]. Hierarchies [413]. High [356, 220, 219, 112, 247, 294]. High-Contrast [356, 220, 247]. High-Dimensional [219]. Higher [3, 133]. Highway [31]. Hilliard [104, 272]. HIV [59]. Holling [80]. Homogeneous [404]. Homogenization [167, 233, 247]. Homogenized [147]. Honey [40]. Honeycomb [393]. Hopf [303]. Hospitals [340]. Host [378]. Howie [191]. Hückel [92].

Human [84]. Hunger [118]. Hybrid [117]. Hydrodynamic [364]. Hydrodynamics [145, 390]. Hyperbolic [108, 271]. Hypergraph [323]. Hypersingular [101]. Hypocoercivity [55]. Hypothesis [410]. Hysteresis [105]. Hysteretic [388].

Ice [188, 302, 294]. Ice-Fishing [294]. Ideal [155]. Identifies [372]. Ignition [122]. II [133, 3, 332]. III [164]. Imaging [257, 401, 265]. Immersed [297]. Immune [336, 408]. Immunity [375]. Impact [68, 173]. Impedance [258, 5, 228, 159, 196, 25]. Impenetrable [393]. Implementation [152]. Implications [331]. Implicit [331]. Impulsive [26, 81]. In-Plane [222]. Inactivation [120]. Incidence [344, 39]. Including [259, 285]. Inclusion [179]. Inclusions [302, 110, 65]. Incompatible [152]. Increase [362]. Induced [135, 165]. Inertia [175]. Infection [98]. Infections [176, 277, 307]. Infectious [277]. Infectivity [206]. Inference [322, 373]. Infinite [70, 275, 351]. Infinitely [87, 341]. Inflated [61]. Infused [381]. Inhomogeneities [388]. Inhomogeneous [190, 114, 200, 353]. Initial [268]. Injection [289]. Insect [192, 152]. Insoluble [232]. Instabilities [135]. Instability [388]. Insulators [396]. Insulin [238]. Intake [354]. Integral [301, 208, 369, 328]. Integrals [221]. Integrating [288]. Integro [290]. Integro-Differential [290]. Interacting [263, 309, 304, 24, 15]. Interaction [28, 406, 96, 408]. Interactions [118, 291, 110, 352]. Interdigitated [38]. Interface [162, 282, 21, 75, 272, 285]. Interfaces [248, 388]. Interplay [33]. Intervention [344]. Intimate [194]. Invasion [173, 44, 320]. Inverse [202, 67, 42, 379, 2, 181, 121, 18, 391, 3, 208, 265, 114, 142, 231, 286, 397, 347, 383, 270, 412, 346, 29, 115, 341, 196, 217, 103, 133].

Inversion [162]. Inviscid [187]. Involving [105]. Ion [266, 154]. Ions [170]. Iontophoretic [85]. Isentropic [111]. Isotherm [197]. Isothermal [95].

Jams [367]. Jets [122]. Joint [137]. Jones [282]. Joseph [282, 282]. Jump [101, 324]. Jump-Diffusions [324]. Junction [17, 51].

Keller [40, 14]. Kernel [119]. Kernels [101]. Kerr [353]. Key [180]. Kinematic [371]. Kinetic [271, 13, 61, 216]. Kirchhoff [70]. Kolmogorov [81]. Koshland [52]. KP [332]. KP-II [332]. Kuramoto [295]. Kuramoto-Oscillator [295].

Laminar [122]. Landscape [316].

Langevin [55]. Laplace [299]. Larché [99]. Large [236, 18, 324, 35]. Large-Population [236]. Lattice [393, 332, 406]. Lattices [396, 207]. Lavrent’ev [208]. Law [297, 134, 35]. Laws [210]. Layer [333, 27, 163, 357]. Layered [247, 312]. Layers [210]. Learning [306]. Length [165]. Level [162]. Lévy [54, 410, 24]. Lexicographic [180]. Liable [241]. Life [176]. Life-Saving [176]. Life-Threatening [176]. Lifshitz [143]. Lift [146]. Lift-Off [146]. Light [376, 352]. Like [150]. Limit [83, 145, 236, 187, 364, 272]. Limits [24, 246]. Line [21]. Linear [138, 187, 208, 228, 252, 134, 62, 196]. Linearized [18]. Linker [285]. Liquid [93, 135, 8, 316]. Lithium [266, 154]. Lithium-Ion [266, 154]. Loaded [355]. Local [30, 260, 113, 184]. Localization [254, 166, 261, 373]. Localized [138, 159]. Locally [234, 23]. Locating [197]. Logistic [127]. Long [330, 6, 372, 256, 39, 50, 406]. Long-Range [330, 406]. Long-Term [372, 39]. Longitudinal [38, 4]. Looping [160]. Loss [157]. Lotka [193]. Lottery [230]. Love [70].

Macro [83]. Macroscopic [128, 31, 414, 246]. Magnetic [233, 1, 91, 351]. Magnetohydrodynamic [100]. Malaria [375]. Many [87, 164, 341]. Many-to-One [164]. Map [117]. Mapping [299]. Maps [363, 185, 62]. Marangoni [232, 8]. Marangoni-Driven [8]. Markov [389, 387]. Mass [87]. Mass-Action [87]. Master [364]. Matched [357]. Matching [293, 215]. Material [355]. Materials [245, 247]. Mathematical [189, 67, 32, 176, 27, 94, 150, 362, 178, 89]. Mathematics [36]. Matrices [267]. Matrix [36, 312]. Mattress [4]. Maximization [321]. Maximizing [209]. Maxwell [345, 353, 312]. Mean [53, 100, 54, 22, 348, 388, 323, 129, 35]. Mean-Field [53, 100, 348, 388, 129]. Mean-Return-Time [22]. Measurement [385, 379, 397]. Measurements [385, 250]. Measures [239, 342, 261, 322]. Mechanics [243, 338, 212]. Mechanism [343]. Mechanisms [394, 60]. Media [330, 392, 356, 190, 398, 114, 23, 212, 312]. Mediated [278, 354]. Medium [2, 181, 283, 270, 200, 353]. Meinhardt [138]. Melanoma [43]. Membrane [279, 110, 69, 285, 19]. Memory [351]. Meshkov [165]. Meso [246]. Metabolic [238]. Metastable [284]. Method [308, 213, 293, 271, 403, 357, 250, 346, 71, 151, 103]. Methods [164, 13, 237]. Metric [49]. Metropolized [386]. Micro [83, 246]. Micro-Macro [83]. Micro-Meso [246]. Microalgae [37]. Microbial [335]. Microcavities [89]. Microorganisms [20]. Microscale [99]. Microscopic [63]. Microsilica [47]. Migration [141, 302, 246]. Minimal [385, 326]. Minimization [73]. Minimizers [104]. Minimizing [344, 405]. Minimum [304]. Minnaert [245]. Mirage [240]. Mitral [243]. Mixed [219, 59]. Mixed-Dimensional [219]. Mixed-Effects [59]. Mixing [153]. Mixotrophic [106].

Modal [119]. Mode [38, 346]. Model [68, 255, 40, 125, 207, 160, 161, 141, 189, 407, 14, 314, 230, 83, 238, 137, 258, 43, 278, 193, 105, 108, 340, 176, 27, 38, 138, 197, 74, 56, 388, 80, 82, 249, 61, 75, 112, 77, 136, 147, 334, 128, 59, 28, 44, 395, 195, 327, 252, 49, 374, 149, 39, 274, 289, 362, 178, 69, 273, 84, 411, 31, 336, 414, 280, 307, 227, 154, 366, 117, 319, 51, 339, 285, 288, 182, 169, 81, 328, 408, 244]. Modeled [257]. Modeling [40, 144, 259, 11, 63, 387, 47, 34, 266, 335, 37, 126, 345, 320, 375, 99, 290, 212, 140, 152]. Models [100, 144, 66, 30, 310, 26, 317, 206, 132, 323, 20, 219, 236, 394, 1, 91, 398, 143, 390, 322, 168, 264, 127, 86, 177, 203, 225, 218, 311, 60, 184]. Modified [250, 195]. Modularity [321]. Modulated [93, 347]. Molecular [390]. Molecular-Theory-Based [390]. Molecules [113]. Moment [13, 216]. Moments [271]. Momentum [235]. Monolayer [361]. Monotone [317]. Monotonicity [267, 2, 264, 181]. Monoxide [47]. Monte [386]. Morphogenesis [161]. Morphomechanical [160]. Mortal [120]. Mosquito [320, 152]. Mosquito-Borne [320]. Mosquitoes [44]. Most [410]. Motion [263]. Motions [215, 107]. Mouth [307]. Movement [229, 139, 212, 117]. Moving [331, 197, 366, 373]. MRI [333]. Mucus [163]. Multi [346, 351]. Multi-Dimensional [351]. Multi-Frequency [346]. Multifrequency [226, 409]. Multilane [414]. Multilayer [239, 342, 141]. Multilayered [392]. Multiple [56, 106, 102, 96]. Multiple-Scale [56]. Multiply [369]. Multiscale [259, 183, 409, 99, 288]. Multistability [269]. Multistate [322]. Myocardial [288]. Myosin [273]. Myosin-Driven [273]. Myxomatosis [149].

Nabarro [112]. Narrow [94, 298]. Navier [297]. Navigation [84]. Near [53, 21, 296, 250, 327, 198]. Near-Collision [327]. Near-Exact [296]. Necessary [144]. Needle [287]. Negative [61]. Nematic [135, 76, 316]. Nematics [390]. Nernst [144, 195, 86, 345]. Network [194, 44]. Networks [79, 360, 239, 342, 167, 389, 63, 348, 132, 372, 306, 194, 173, 295, 185, 153, 337, 58, 35, 402, 115, 405]. Neumann [73]. Neural [203, 225, 35, 115]. Neurodegenerative [337]. Neuronal [144]. Neutral [65]. Neutron [164]. Newtonian [343]. Next [189]. Next-Generation [189]. No [303]. No-Contrast [303]. Noise [350, 156, 387, 295, 380]. Noisy [241, 327]. Non [343]. Non-Newtonian [343]. Noncompliant [411]. Nonconvex [357]. Nondegenerate [360]. Nondilute [233]. Nonelliptical [65]. Nonenergetic [127]. Nonlinear [396, 257, 175, 64, 78, 329, 368, 108, 399, 77, 187, 208, 124, 398, 265, 69, 402, 353, 268, 339, 217, 130]. Nonlinearities [33]. Nonlocal [125, 83, 20, 249, 307]. Nonnegative [267]. Nonradiating [190]. Nonsmooth [358]. Nonstandard [183]. Nonstationary [230]. Nonuniversality [198]. Nonzero [73, 48]. Normal [363]. Novel [267, 238, 204, 246]. Number [57, 90, 385]. Numbers [35]. Numerical [263, 67, 331, 208, 260, 153, 290, 89, 182]. Nutrient [183, 227]. Nutrient-Phytoplankton-Zooplankton [227]. Nutrients [352].

Object [326]. Observation [123, 391]. Obstacles [297, 393, 270]. Odd [368]. ODE [69]. Off [146]. Ogden [96]. Oil [141]. One [144, 164, 269, 356, 271, 69, 62, 200, 268]. One-Dimensional [269, 356, 271, 69, 62, 200, 268]. Only [385]. Operator [202, 231]. Operators [159, 131]. Opinion [348, 211]. Optical [401, 89]. Optimal [207, 344, 318, 314, 10, 370, 5, 382, 171, 335, 127, 153, 84, 29, 169]. Optimization [162, 172, 64, 127, 153].

Optimizing [241]. Orb [222]. Orb-Web [222]. Orbits [317]. Order [269, 132, 283, 187, 3, 133, 149, 338]. Ordinary [34]. Orthogonal [93]. Oscillations [380, 289, 402]. Oscillator [295]. Oscillators [22]. Oscillatory [79, 69]. Other [331]. Ovarian [255]. Overlay [306]. Overwashed [157]. Oxidation [102]. Oxygen [331, 47].

Packed [370]. Pairwise [291]. Parameter [137, 196]. Parameters [180, 338]. Parametric [284]. Paraxial [330]. Parcel [34]. Partial [213, 22]. Partially [298]. Particle [392, 158, 47, 236, 266, 28]. Particle-Size [266]. Particles [309, 233, 304, 24, 381]. Partitioned [298]. Partitions [386]. Passage [120]. Passive [379, 156]. Pasta [332]. Patch [264, 355, 381]. Pathogen [325]. Pattern [314, 214, 121, 402]. Patterns [349, 138, 286, 409, 380]. PB [395]. PB-Steric [395]. PDE [69, 375, 117, 60]. PDEs [220, 236]. Pedestrian [305]. Peierls [112]. Pelagic [352]. Penetrable [303]. Percolation [58]. Perfectly [357]. Perfusion [288]. Period [62]. Periodic [167, 317, 234, 94, 264, 23, 151, 404, 131]. Periodical [252]. Periodicity [155]. Permanent [86]. Permeability [17]. Persistence [193, 117]. Perturbation [295]. Perturbed [365, 267, 234]. Pest [378, 81]. Pest-Host [378]. Phase [309, 22, 407, 14, 385, 72, 388, 73, 28, 380, 99, 212, 115]. Phase-Field [28, 212]. Phase-Only [385]. Phaseless [103]. Phenomena [116]. Phenomenon [127]. Phenotypic [140]. Phonon [248]. Photoacoustic [401, 265, 374]. Physical [33]. Physiology [201]. Phytoplankton [252, 227]. Picard [412]. Piecewise [387, 252, 62]. Piezoelectric [351]. Planar [22]. Planck [144, 195, 86, 345]. Plane [123, 222, 298, 341, 96]. Plant [354, 183].

Plasmon [159]. Plasmonics [401]. Plasmons [166]. Plasticity [203, 225]. Plate [70, 379]. Platooning [31]. Plesset [170]. PML [403]. Point [276, 268]. Point-Source [268]. Points [393]. Poisson [144, 92, 395, 195, 86]. Polarizability [326]. Polarization [72, 93]. Polarization-Modulated [93]. Policies [171]. Polluted [244]. Polyelectrolyte [377]. Polygons [76, 48]. Polyhedral [121]. Polynomial [400]. Polytopes [215]. Population [68, 255, 267, 186, 82, 236, 127, 71, 152]. Populations [244]. Poromechanics [288]. Porous [23, 212]. Posedness [70, 219]. Positive [87]. Posteriori [259]. Potato [146]. Potential [18, 86]. Potentials [333]. Pouch [154]. POVMs [281]. Power [175, 134]. Pre [362]. Pre-Deterioration [362]. Predator [118, 80, 274, 280]. Predators [80]. Predicting [146]. Prescribed [9]. Presence [96]. Preserves [360]. Preserving [384]. Pressure [213]. Preventing [378]. Prey [118, 80, 222, 274, 280]. Principal [46, 264]. Probabilistic [148]. Problem [123, 202, 6, 67, 331, 42, 379, 387, 45, 301, 382, 18, 73, 208, 174, 114, 231, 397, 347, 127, 346, 89, 130, 294, 196, 217, 355]. Problems [70, 331, 219, 403, 391, 3, 133, 265, 286, 260, 270, 115, 341]. Procedures [176]. Process [259, 172, 387, 382]. Processes [359]. Producers [352]. Production [27, 328]. Profile [21]. Profiles [57, 90]. Progression [362]. Projection [46]. Proliferation [60]. Promoter [394, 322]. proofs [267]. Propagation [330, 186, 329, 26, 403, 398, 129, 168, 252, 337, 117]. Properties [240, 401, 55, 361, 336]. Proteins [19]. Protocols [32].

Quadrature [271, 13]. Quadrature-Based [13]. Qualitative [263]. Quantification [256]. Quantum [239, 342]. Quarter [123].

Quarter-Plane [123]. Quasi [301, 298, 218]. Quasi-Bound [298]. Quasi-Resonances [301]. Quasi-Stationarity [218]. Queueing [171, 405]. Queuing [221]. Queuing-Type [221].

Radial [111]. Radially [339]. Radiating [296]. Radiation [150, 237]. Radiative [108, 124]. Radius [145]. Radon [358]. Random [330, 239, 342, 142, 223, 402, 184]. Random-Walk [239, 342]. Randomly [200]. Range [330, 406]. Rascle [83]. Rate [284]. Rates [354, 39]. Ratios [404]. Rayleigh [170, 165]. Reactants [102]. Reaction [229, 79, 360, 125, 11, 314, 389, 95, 310, 26, 47, 372, 236, 313, 364, 102, 411, 130, 268, 311, 205, 182, 169]. Reaction-Diffusion [314, 236, 313, 364, 411, 268, 311, 169]. Reaction-Time [310]. Reactions [79, 360]. Reactive [122, 147]. Reactor [366]. Reactors [172]. Real [293]. Realizations [400]. Recombination [386]. Reconstructing [248, 287, 412]. Reconstruction [213, 385, 269, 401, 226, 358, 124, 409]. Reconstructions [162]. Recovery [283]. Recruitment [192]. Rectangular [316]. Reduced [76]. Reduction [382, 380, 154]. Reflected [324, 157]. Refugee [212]. Regeneration [278]. Regular [101, 48]. Regularity [412]. Regularization [412, 9]. Regulations [238]. Relation [194, 143, 353]. Relations [101, 281]. Relay [319]. Relevant [174]. Reliable [100, 306]. Renewal [392]. Representation [333, 279]. Representations [199]. Reproduction [57, 90, 404]. Residence [261]. Resolution [358]. Resonance [284, 1, 91, 94]. Resonances [356, 301, 245, 242, 159]. Resonant [188, 298]. Resonators [308, 77]. Resource [354]. Resource-Mediated [354]. Respect [317]. Response [15, 280, 408]. Resting [329]. Results [204].

Return [22]. Reversal [86]. Reversible [79, 87, 12, 400]. Revisited [303]. Rewarding [410]. Richtmyer [165]. Right [303]. Right-Angled [303]. Rigid [145, 215, 113]. Rigid-Body [215]. Rigorous [246]. Ring [367]. Riots [180]. River [244]. RNA [61]. Road [367]. Robotic [24]. Robustness [315]. Rolling [253]. Root [300, 183]. Roots [183, 221]. Rotational [19]. Route [364].

Saffman [282, 343]. Sampling [386, 250]. Saturated [106]. Saturation [384]. Savanna [334]. Saving [176]. Scaffold [278]. Scalar [119, 220, 332]. Scale [56, 165]. Scale-Dependent [165]. Scaling [183, 225]. Scattered [226]. Scatterers [121, 276]. Scattering [70, 42, 2, 181, 121, 234, 191, 398, 142, 383, 412, 103]. Scheme [9]. Schemes [384]. Schrödinger [18, 191]. Schwarz [369]. Sea [302]. Search [161, 287]. Search-and-Capture [161]. Searchers [120]. Searching [300, 80]. Seasonal [334]. Second [132, 41]. Second-Order [132]. Segel [40, 14]. Segregation [311]. SEIRS [317]. Selection [343, 10, 16]. Self [158, 165]. Self-Driven [158]. Self-Similar [165]. Semi [355]. Semi-space [355]. Sensitivity [389]. Sensors [148, 374]. Separation [309, 99]. Sequencing [61]. Series [134]. Set [162]. Set-Based [162]. Sexual [186]. Shallow [368, 134]. Shallow-Water [368]. Shannon [281]. Shape [162, 172, 215]. Shapes [65]. Sharp [162, 75, 272]. Sharp-Interface [162, 75]. Shaw [16]. Shear [237]. Sheaves [211]. Shelf [188]. Shells [304]. Shelter [340]. Shift [117]. Shifting [193]. Sigmoid [280]. Sign [40]. Sign-Changing [40]. Signal [385, 333, 403, 246]. Signal-Propagation [403]. Signal-Triggered [246]. Signaling [144]. Signals [362]. Significance [149]. Silicon [47]. Similar [413, 165]. Similarity [413]. Simple [148]. Simulate [176].

Simulating [75]. Simulation [159, 99]. Simulations [182]. Simultaneous [401]. Single [392, 367, 121, 61, 346, 405]. Single-Cell [61]. Single-Particle [392]. Single-Sink [405]. Single-Source [405]. Singularly [365]. Sink [405]. SIQR [319]. SIR [314, 411, 218]. SIRS [39]. SIS [218]. Size [98, 266, 127]. Slab [108]. Sliding [107]. Slit [298]. Slits [94]. Slope [8]. Slow [255, 138, 302]. Slow-Fast [255]. Slyozov [143, 143]. Smectic [93]. Snacks [146]. Social [192]. Søderberg [197]. Soft [33, 245]. Soil [300]. Solid [248, 75, 366]. Solid-State [75]. Solution [208, 97, 316]. Solutions [6, 14, 193, 105, 315, 143, 9, 275, 134, 177]. Solvability [34]. Solvent [170]. Some [361, 204, 177]. Source [226, 114, 142, 231, 286, 347, 346, 268, 405, 25]. Sources [190, 409, 373]. Space [263, 293, 115, 355]. Sparse [314, 46, 226, 183, 409]. Spatial [155, 277, 320, 311]. Spatially [380, 185, 205]. Spatio [227]. Spatio-temporal [227]. Spatiotemporal [244]. Special [372]. Speciation [259]. Species [229, 193, 354, 335, 204]. Species-Dependent [229]. Spectral [202, 267, 8, 292]. Spectrum [131]. Speed [82]. Speeds [168]. sphere [195]. Spherical [304]. Spider [222]. Spines [273]. Spirally [247]. Splitting [360]. Spontaneous [158]. Spot [138]. Spots [294]. Spread [193, 277]. Spreading [57, 90, 378]. Spurious [301]. Square [54, 332]. Stability [310, 138, 13, 249, 114, 231, 412, 414, 280, 373, 7, 184, 351, 74]. Stabilization [106]. Stable [177]. Standing [313]. State [137, 1, 91, 75, 177]. State-Parameter [137]. States [254, 87, 389, 304, 298, 41, 44]. Static [174, 102]. Stationarity [322, 218]. Status [336]. Steady [297, 87, 389, 1, 91, 177]. Steady-State [177]. Steigmann [96]. Stekloff [42].

Stents [189]. Step [150]. Step-Like [150]. Steric [395]. Sterile [152]. Stick [322]. Stick-Breaking [322]. Stieltjes [287]. Stochastic [255, 262, 22, 164, 291, 158, 382, 236, 112, 286, 322, 39, 71, 273, 84, 203, 225, 371]. Stoichiometric [56]. Stokes [297, 101]. Stokesian [145]. Straight [96]. Strains [125]. Strategies [344, 410, 24]. Stream [82]. Stream-Population [82]. String [329, 310]. Strong [139, 104]. Structure [118, 30, 251, 50, 62, 58, 107]. Structured [17, 399, 375, 281]. Structures [213, 292, 174, 242, 364]. Study [67, 76, 298]. Subdiffusion [229, 391]. Subject [380]. Substrate [329]. Subwavelength [308, 356]. Succession [68]. Sugimoto [77]. Summation [151]. Supersaturated [143]. Supervised [318]. Suppresses [413]. Suppression [152]. Surface [166, 27, 150, 159, 294, 96, 355]. Surfaces [101, 19]. Surfactant [232, 126]. Surgical [176]. Surprising [123]. Suspended [361]. Suspension [233]. Swarm [24]. Switch [52]. Switching [17, 394]. Symmetric [339]. Symmetries [113]. Symmetry [30, 326, 221, 50]. Synaptic [203, 225]. Synchronization [175, 295]. System [11, 175, 378, 34, 315, 332, 99, 406, 408, 351]. Systematic [78]. Systems [308, 384, 87, 309, 12, 400, 24, 158, 215, 399, 171, 241, 221, 404, 205].

Tails [198]. Taylor [343, 165]. Technique [152]. temporal [227]. Temporally [109]. Tension [294]. Tensor [390]. Tensors [326, 113]. Term [372, 39]. Terminal [391]. Their [396, 6, 287, 372, 298]. Theorems [145]. Theoretic [66]. Theoretical [253]. Theory [235, 92, 390, 94, 383, 270, 69, 157]. Therapy [43, 408]. Thermal [248, 15, 361]. Thermally [15]. Thermistor [197]. Thermo [374]. Thermocapillary [4]. Thermoelectric [361]. Theta [207]. Thin [209, 72, 8]. Third [187]. Third-Order [187]. Threatening [176]. Three [166, 144, 67, 138, 75, 178]. Three-Dimensional [67, 138]. Three-Tiered [178]. Tiered [178]. Tight [308]. Tight-Binding [308]. Tilt [292]. Time [162, 146, 365, 6, 155, 22, 261, 310, 121, 234, 394, 391, 295, 264, 383, 260, 97, 218, 272, 117, 404]. Time- [383]. Time-Dependent [394]. Time-Domain [162]. Time-Fractional [272]. Time-Harmonic [121, 234, 260]. Time-Periodic [264, 404]. Times [120, 337]. Tissue [339]. Tomography [258, 5, 88, 25]. Topological [396]. Topologies [372]. Topology [194]. Torsion [48]. Torsional [160]. Torus [88, 185]. Total [344, 64, 127, 103]. Total-Field [103]. Toxicants [244]. tQSSA [52]. Trace [42, 374]. Tracing [299, 296]. Traction [33]. Traffic [63, 105, 132, 367, 249, 128, 31, 414]. Transdermal [85]. Transfer [108, 124]. Transform [358]. Transient [56, 256, 130, 268]. Transistors [116]. Transition [284, 72, 157]. Transition-Loss [157]. Transitions [14]. Translation [372, 131]. Transmission [240, 340, 248, 45, 301, 149]. Transmitted [157]. Transplant [171]. Transport [349, 318, 10, 164, 24, 116, 248, 398, 265, 163, 153, 23, 29, 200]. Trapping [141]. Traveling [122, 82]. Travelling [74]. Tridimensional [349]. Triggered [246]. Triple [51]. Truncated [242]. Truncating [134]. Truss [174]. Tube [213]. Tubes [209]. Tumor [137, 43, 168, 362, 169, 408]. Tumor-Immune [408]. Tumors [140]. Turbines [253]. Turbulent [126]. Two [407, 267, 193, 370, 354, 38, 76, 313, 73, 28, 44, 335, 374, 159, 204, 89, 196]. Two-Dimensional [370, 38, 76, 89, 196]. Two-Phase [407, 73, 28]. Two-Species [193, 335]. Two-Way [374]. Type [83, 221, 351].

Ulam [332]. Ultrasound [257]. Unbounded [2, 181, 313]. Uncovering [407]. Understanding [168]. Unequal [86]. Unified [396, 362]. Uniform [52, 7]. Unimodal [266]. Unique [34]. Uniqueness [385, 290, 373, 341]. United [44]. Unknown [283, 391]. Unreliability [410]. Upscaling [78, 23]. Uptake [78, 183]. Uranium [102]. Using [180, 297, 40, 259, 64, 333, 401, 367, 326, 364].

Vacuum [6]. Validity [308]. Value [331, 174]. Values [287]. Valve [243]. Vanishing [145]. Vapor [102]. Variable [46]. Variation [64]. Variational [72, 237, 275]. Varying [206, 173, 165]. Vascularized [140]. Vector [149]. Vectorial [220]. Vehicle [367]. Vehicles [249]. Vehicular [63]. Versus [383]. Vertical [368]. Vertices [313]. Vesicles [273]. Vessels [288]. Via [261, 10, 46, 56, 129, 296, 322, 16, 86, 159, 41]. Vibration [222]. Vibrations [188]. Viewpoint [292]. Viral [125]. Visco [374]. Visco-Thermo-Acoustic [374]. Viscoelasticity [3, 196, 133]. Viscosity [368]. Viscous [297, 232, 233]. Visual [49]. Volterra [193]. Volume [384, 14, 9]. Vortex [93, 252, 48]. Voter [30].

Wagner [143]. Walk [239, 342]. Walks [54, 223, 402]. Walls [53]. Warm [34]. Wasserstein [350]. Water [368, 354, 102, 134]. Wave [202, 119, 330, 329, 158, 398, 286, 150, 252]. Waveform [162]. Waveguide [131]. Waveguides [346]. Wavenumber [293, 18]. Waves [368, 74, 82, 313, 190, 347, 380, 276, 320, 157, 353, 341]. Way [374]. Weakly [87, 12, 400, 69]. Wealth [198]. Web [222]. Wedge [303, 276]. Weighted [174]. Well [70, 282, 219]. Well-Posedness [70, 219]. Westervelt [257]. Wetting [21]. Whelan [191]. White [328]. Wiener [303]. Wine [290]. Winter [40]. Within [78]. Without [295]. Wobbling [305]. Wolbachia [320, 152]. Wound [247].

Year [302].

Zero [400, 61]. Zero-Inflated [61]. Zooplankton [227].

References

Jiang:2011:ASS

[1] Yu Jiang, Hiroshi Fujiwara, and Gen Nakamura. Approximate steady state models for magnetic resonance elastography. SIAM Journal on Applied Mathematics, 71(6):1965–1989, ???? 2011. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://epubs.

                                             siam.org/siap/resource/1/smjmap/

v71/i6/p1965_s1. See erratum [91].

Griesmaier:2018:MIM

[2] Roland Griesmaier and Bastian Harrach. Monotonicity in inverse medium scattering on unbounded domains. SIAM Journal on Applied Mathematics, 78(5):2533–2557, ???? 2018. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [181].

Kirsch:2019:IPA

[3] Andreas Kirsch and Andreas Rieder. Inverse problems for abstract evolution equations II: Higher order differentiability for viscoelasticity. SIAM Journal on Applied Mathematics, 79 (6):2639–2662, ???? 2019. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [133].

Yariv:2020:LTF

[4] Ehud Yariv and Darren Crowdy. Longitudinal thermocapillary flow over a dense bubble mattress. SIAM Journal on Applied Mathematics, 80(1):1–19, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garde:2020:ODD

[5] Henrik Garde and Nuutti Hyvönen. Optimal depth-dependent distinguishability bounds for electrical impedance tomography in arbitrary dimension. SIAM Journal on Applied Mathematics, 80 (1):20–43, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Camassa:2020:VDB

[6] Roberto Camassa, Gregorio Falqui, Giovanni Ortenzi, Marco Pedroni, and Giuseppe Pitton. On the “Vacuum” dam-break problem: Exact solutions and their long time asymptotics. SIAM Journal on Applied Mathematics, 80 (1):44–70, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yorkston:2020:DSE

[7] Adam A. Yorkston, Mark G. Blyth, and Emilian I. Parau. The deformation and stability of an elastic cell in a uniform flow. SIAM Journal on Applied Mathematics, 80(1):71–94, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ghazaryan:2020:SAF

[8] Anna Ghazaryan, Stephane Lafortune, and Vahagn Manukian. Spectral analysis of fronts in a Marangoni-driven thin liquid film flow down a slope. SIAM Journal on Applied Mathematics, 80(1):95–118, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Negron-Marrero:2020:CRS

[9] Pablo V. Negrón-Marrero and Jeyabal Sivaloganathan. On the convergence of a regularization scheme for approximating cavitation solutions with prescribed cavity volume. SIAM Journal on Applied Mathematics, 80(1):119–141, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chow:2020:ESO

[10] Shui nee Chow, Wuchen Li, Jun Lu, and Haomin Zhou. Equilibrium selection via optimal transport. SIAM Journal on Applied Mathematics, 80 (1):142–159, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Burger:2020:RCD

[11] Martin Burger, Patricia Friele, and Jan-Frederik Pietschmann. On a reaction–cross-diffusion system modeling the growth of glioblastoma. SIAM Journal on Applied Mathematics, 80 (1):160–182, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Craciun:2020:ECC

[12] Gheorghe Craciun, Jiaxin Jin, and Polly Y. Yu. An efficient characterization of complex-balanced, detailedbalanced, and weakly reversible systems. SIAM Journal on Applied Mathematics, 80(1):183–205, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huang:2020:SAQ

[13] Qian Huang, Shuiqing Li, and Wen-An Yong. Stability analysis of quadraturebased moment methods for kinetic equations. SIAM Journal on Applied Mathematics, 80(1):206–231, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carrillo:2020:PTB

[14] Jose A. Carrillo, Xinfu Chen, Qi Wang, Zhian Wang, and Lu Zhang. Phase transitions and bump solutions of the Keller–Segel model with volume exclusion. SIAM Journal on Applied Mathematics, 80(1):232–261, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hermanns:2020:HTR

[15] Miguel Hermanns and Santiago Ibáñez. Harmonic thermal response of thermally interacting geothermal boreholes. SIAM Journal on Applied Mathematics, 80(1):262–288, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lustri:2020:SHS

[16] Christopher J. Lustri, Christopher C. Green, and Scott W. McCue. Selection of a Hele–Shaw bubble via exponential asymptotics. SIAM Journal on Applied Mathematics, 80(1):289–311, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bressloff:2020:EPG

[17] Paul C. Bressloff, Sean D. Lawley, and Patrick Murphy. Effective permeability of a gap junction with age-structured switching. SIAM Journal on Applied Mathematics, 80(1):312–337, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isakov:2020:LIS

[18] Victor Isakov, Shuai Lu, and Boxi Xu. Linearized inverse Schrödinger potential problem at a large wavenumber. SIAM Journal on Applied Mathematics, 80(1):338–358, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhou:2020:CDR

[19] Yongcheng Zhou. On curvature driven rotational diffusion of proteins on membrane surfaces. SIAM Journal on Applied Mathematics, 80(1):359–381, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hillen:2020:NAM

[20] Thomas Hillen and Andreas Buttenschön. Nonlocal adhesion models for microorganisms on bounded domains. SIAM Journal on Applied Mathematics, 80(1):382–401, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cui:2020:IPN

[21] Hanwen Cui and Weiqing Ren. Interface profile near the contact line in electro-wetting on dielectric. SIAM Journal on Applied Mathematics, 80 (1):402–421, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cao:2020:PDE

[22] Alexander Cao, Benjamin Lindner, and Peter J. Thomas. A partial differential equation for the mean-return-time phase of planar stochastic oscillators. SIAM Journal on Applied Mathematics, 80(1):422–447, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Muntean:2020:CTL

[23] Adrian Muntean and Christos Nikolopoulos. Colloidal transport in locally periodic evolving porous media — an upscaling exercise. SIAM Journal on Applied Mathematics, 80(1):448–475, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Estrada-Rodriguez:2020:IPL

[24] Gissell Estrada-Rodriguez and Heiko Gimperlein. Interacting particles with Lévy strategies: Limits of transport equations for swarm robotic systems. SIAM Journal on Applied Mathematics, 80(1):476–498, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2020:SCE

[25] Tingting Zhang, Geuk Young Jang, Tong In Oh, Kyung Woon Jeung, Hun Wi, and Eung Je Woo. Source consistency electrical impedance tomography. SIAM Journal on Applied Mathematics, 80(1):499–520, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fazly:2020:API

[26] Mostafa Fazly, Mark Lewis, and Hao Wang. Analysis of propagation for impulsive reaction–diffusion models. SIAM Journal on Applied Mathematics, 80 (1):521–542, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gandolfi:2020:BLE

[27] Alberto Gandolfi, Mimmo Iannelli, and Gabriela Marinoschi. The basal layer of the epidermis: a mathematical model for cell production under a surface density constraint. SIAM Journal on Applied Mathematics, 80(1):543–571, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:2020:EDP

[28] Xiang Li, Qiang Du, and Xiao-Ping Wang. Energy decaying phase-field model for fluid-particle interaction in two-phase flow. SIAM Journal on Applied Mathematics, 80(1):572–598, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stuart:2020:IOT

[29] Andrew M. Stuart and Marie-Therese Wolfram. Inverse optimal transport. SIAM Journal on Applied Mathematics, 80(1):599–619, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chodrow:2020:LSG

[30] Philip S. Chodrow and Peter J. Mucha. Local symmetry and global structure in adaptive voter models. SIAM Journal on Applied Mathematics, 80 (1):620–638, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Piacentini:2020:MMP

[31] Giulia Piacentini, Paola Goatin, and Antonella Ferrara. A macroscopic model for platooning in highway traffic. SIAM Journal on Applied Mathematics, 80 (1):639–656, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dalwadi:2020:MFD

[32] Mohit P. Dalwadi, Sarah L. Waters, Helen M. Byrne, and Ian J. Hewitt. A mathematical framework for developing freezing protocols in the cryopreservation of cells. SIAM Journal on Applied Mathematics, 80(2):657–689, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ciarletta:2020:EFB

[33] Pasquale Ciarletta, Hui-Hui Dai, and Matteo Taffetani. Elastic fingering of a bonded soft disc in traction: Interplay of geometric and physical nonlinearities. SIAM Journal on Applied Mathematics, 80(2):690–705, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hanke:2020:USS

[34] Martin Hanke and Nikolas Porz. Unique solvability of a system of ordinary differential equations modeling a warm cloud parcel. SIAM Journal on Applied Mathematics, 80(2):706–724, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Sirignano:2020:MFA

[35] Justin Sirignano and Konstantinos Spiliopoulos. Mean field analysis of neural networks: a law of large numbers.

SIAM Journal on Applied Mathematics, 80(2):725–752, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zabarankin:2020:MDE

[36] Michael Zabarankin and Bogdan Grechuk. Matrix difference equations in applied mathematics. SIAM Journal on Applied Mathematics, 80(2):753–771, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Martinez:2020:MAA

[37] Carlos Martinez, Andrés Ávila, Francis Mairet, Leslie Meier, and David Jeison. Modeling and analysis of an absorption column connected to a microalgae culture. SIAM Journal on Applied Mathematics, 80(2):772–791, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gaudiello:2020:TDE

[38] Antonio Gaudiello and Michel Lenczner. A two-dimensional electrostatic model of interdigitated comb drive in longitudinal mode. SIAM Journal on Applied Mathematics, 80(2):792–813, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nguyen:2020:LTA

[39] Dang Hai Nguyen, George Yin, and Chao Zhu. Long-term analysis of a stochastic SIRS model with general incidence rates. SIAM Journal on Applied Mathematics, 80(2):814–838, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bastiaansen:2020:MHB

[40] Robbin Bastiaansen, Arjen Doelman, Frank van Langevelde, and Vivi Rottschäfer. Modeling honey bee colonies in winter using a Keller–Segel model with a sign-changing chemotactic coefficient. SIAM Journal on Applied Mathematics, 80(2):839–863, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yuan:2020:EBS

[41] Lijun Yuan and Ya Yan Lu. Excitation of bound states in the continuum via second harmonic generations. SIAM Journal on Applied Mathematics, 80(2):864–880, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cogar:2020:ATC

[42] Samuel Cogar. Analysis of a trace class Stekloff eigenvalue problem arising in inverse scattering. SIAM Journal on Applied Mathematics, 80(2):881–905, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dickman:2020:TCE

[43] Lauren R. Dickman, Evan Milliken, and Yang Kuang. Tumor control, elimination, and escape through a compartmental model of dendritic cell therapy for melanoma. SIAM Journal on Applied Mathematics, 80(2):906–928, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:2020:NMT

[44] Zuhan Liu, Canrong Tian, and Shigui Ruan. On a network model of two competitors with applications to the invasion and competition of Aedes Albopictus and Aedes Aegypti mosquitoes in the United States. SIAM Journal on Applied Mathematics, 80(2):929–950, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Helsing:2020:ECC

[45] Johan Helsing and Anders Karlsson. An extended charge–current formulation of the electromagnetic transmission problem. SIAM Journal on Applied Mathematics, 80(2):951–976, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Erichson:2020:SPC

[46] N. Benjamin Erichson, Peng Zheng, Krithika Manohar, Steven L. Brunton, J. Nathan Kutz, and Aleksandr Y. Aravkin. Sparse principal component analysis via variable projection. SIAM Journal on Applied Mathematics, 80 (2):977–1002, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gonzalez-Farina:2020:MMP

[47] Raquel González-Fariña, Andreas Münch, James M. Oliver, and Robert A. Van Gorder. Modeling microsilica particle formation and growth due to the combustion reaction of silicon monoxide with oxygen. SIAM Journal on Applied Mathematics, 80(2):1003–1033, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

delaHoz:2020:EVF

[48] Francisco de la Hoz, Sandeep Kumar, and Luis Vega. On the evolution of the vortex filament equation for regular M-polygons with nonzero torsion. SIAM Journal on Applied Mathematics, 80(2):1034–1056, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Montobbio:2020:MMF

[49] Noemi Montobbio, Alessandro Sarti, and Giovanna Citti. A metric model for the functional architecture of the visual cortex. SIAM Journal on Applied Mathematics, 80(2):1057–1081, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pandurangi:2020:HAS

[50] Shrinidhi S. Pandurangi, Timothy J. Healey, and Nicolas Triantafyllidis. Hidden asymptotic symmetry in a long elastic structure. SIAM Journal on Applied Mathematics, 80(3):1083–1100, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wei:2020:GBT

[51] Chaozhen Wei, Luchan Zhang, Jian Han, David J. Srolovitz, and Yang Xiang. Grain boundary triple junction dynamics: a continuum disconnection model. SIAM Journal on Applied Mathematics, 80(3):1101–1122, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bersani:2020:UAE

[52] A. M. Bersani, A. Borri, A. Milanesi, G. Tomassetti, and P. Vellucci. Uniform asymptotic expansions beyond the tQSSA for the Goldbeter–Koshland switch. SIAM Journal on Applied Mathematics, 80(3):1123–1152, ???? 2020.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Aurell:2020:BNW

[53] Alexander Aurell and Boualem Djehiche. Behavior near walls in the mean-field approach to crowd dynamics. SIAM Journal on Applied Mathematics, 80(3):1153–1174, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Borgers:2020:MSD

[54] Christoph Börgers and Claude Greengard. On the mean square displacement in Lévy walks. SIAM Journal on Applied Mathematics, 80(3): 1175–1196, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Leimkuhler:2020:HPA

[55] Benedict Leimkuhler, Matthias Sachs, and Gabriel Stoltz. Hypocoercivity properties of adaptive Langevin dynamics. SIAM Journal on Applied Mathematics, 80(3):1197–1222, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Heggerud:2020:TDS

[56] Christopher M. Heggerud, Hao Wang, and Mark A. Lewis. Transient dynamics of a stoichiometric cyanobacteria model via multiple-scale analysis. SIAM Journal on Applied Mathematics, 80(3):1223–1246, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:2020:APB

[57] Shanshan Chen and Junping Shi. Asymptotic profiles of basic reproduction number for epidemic spreading in heterogeneous environment. SIAM Journal on Applied Mathematics, 80 (3):1247–1271, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [90].

Shang:2020:GCP

[58] Yilun Shang. Generalized K-core percolation in networks with community structure. SIAM Journal on Applied Mathematics, 80(3):1272–1289, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:2020:HME

[59] Yunjeong Lee, Yoon gu Hwang, Hee-Dae Kwon, Jun Yong Choi, and Jeehyun Lee. Hierarchical mixed-effects model for HIV dynamics. SIAM Journal on Applied Mathematics, 80(3): 1290–1306, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xia:2020:PMA

[60] Mingtao Xia, Chris D. Greenman, and Tom Chou. PDE models of adder mechanisms in cellular proliferation. SIAM Journal on Applied Mathematics, 80 (3):1307–1335, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jia:2020:KFZ

[61] Chen Jia. Kinetic foundation of the zero-inflated negative binomial model for single-cell RNA sequencing data. SIAM Journal on Applied Mathematics, 80(3):1336–1355, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Segura:2020:DPA

[62] Juan Segura, Frank M. Hilker, and Daniel Franco. Degenerate period adding bifurcation structure of onedimensional bimodal piecewise linear maps. SIAM Journal on Applied Mathematics, 80(3):1356–1376, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Colombo:2020:MMV

[63] Rinaldo M. Colombo, Helge Holden, and Francesca Marcellini. On the microscopic modeling of vehicular traffic on general networks. SIAM Journal on Applied Mathematics, 80(3): 1377–1391, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cristofari:2020:TVB

[64] Andrea Cristofari, Francesco Rinaldi, and Francesco Tudisco. Total variation based community detection using a nonlinear optimization approach. SIAM Journal on Applied Mathematics, 80(3):1392–1419, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lim:2020:IGS

[65] Mikyoung Lim and Graeme W. Milton. Inclusions of general shapes having constant field inside the core and nonelliptical neutral coated inclusions with anisotropic conductivity. SIAM Journal on Applied Mathematics, 80 (3):1420–1440, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cartee:2020:CTM

[66] Elliot Cartee and Alexander Vladimirsky. Control-theoretic models of environmental crime. SIAM Journal on Applied Mathematics, 80(3):1441–1466, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:2020:MNS

[67] Junqing Chen, Ying Liang, and Jun Zou. Mathematical and numerical study of a three-dimensional inverse eddy current problem. SIAM Journal on Applied Mathematics, 80(3): 1467–1492, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bai:2020:DSM

[68] Zhenguo Bai, Yijun Lou, and Xiao-Qiang Zhao. A delayed succession model with diffusion for the impact of diapause on population growth. SIAM Journal on Applied Mathematics, 80 (3):1493–1519, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Paquin-Lefebvre:2020:WNT

[69] Frédéric Paquin-Lefebvre, Wayne Nagata, and Michael J. Ward. Weakly nonlinear theory for oscillatory dynamics in a one-dimensional PDE–ODE model of membrane dynamics coupled by a bulk diffusion field. SIAM Journal on Applied Mathematics, 80 (3):1520–1545, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bourgeois:2020:WPS

[70] Laurent Bourgeois and Christophe Hazard. On well-posedness of scattering problems in a Kirchhoff–Love infinite plate. SIAM Journal on Applied Mathematics, 80(3):1546–1566, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nguyen:2020:MDC

[71] Dang H. Nguyen and Edouard Strickler. A method to deal with the critical case in stochastic population dynamics. SIAM Journal on Applied Mathematics, 80(3):1567–1589, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Du:2020:VPF

[72] Qiang Du, Ruotai Li, and Lei Zhang. Variational phase field formulations of polarization and phase transition in ferroelectric thin films. SIAM Journal on Applied Mathematics, 80(3): 1590–1606, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kang:2020:MFN

[73] Di Kang, Patrick Choi, and Chiu-Yen Kao. Minimization of the first nonzero eigenvalue problem for two-phase conductors with Neumann boundary conditions. SIAM Journal on Applied Mathematics, 80(4):1607–1628, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Harley:2020:STW

[74] Kristen E. Harley, Peter van Heijster, Robert Marangell, Graeme J. Pettet, Timothy V. Roberts, and Martin Wechselberger. (In)stability of travelling waves in a model of haptotaxis. SIAM Journal on Applied Mathematics, 80 (4):1629–1653, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jiang:2020:SIM

[75] Wei Jiang, Quan Zhao, and Weizhu Bao. Sharp-interface model for simulating solid-state dewetting in three dimensions. SIAM Journal on Applied Mathematics, 80(4):1654–1677, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Han:2020:RSN

[76] Yucen Han, Apala Majumdar, and Lei Zhang. A reduced study for nematic equilibria on two-dimensional polygons. SIAM Journal on Applied Mathematics, 80(4):1678–1703, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Junca:2020:ASM

[77] Stéphane Junca and Bruno Lombard. Analysis of a sugimoto model of nonlinear acoustics in an array of Helmholtz resonators. SIAM Journal on Applied Mathematics, 80(4):1704–1722, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dalwadi:2020:SUN

[78] Mohit P. Dalwadi and John R. King. A systematic upscaling of nonlinear chemical uptake within a biofilm. SIAM Journal on Applied Mathematics, 80 (4):1723–1750, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Banaji:2020:BOC

[79] Murad Banaji. Building oscillatory chemical reaction networks by adding reversible reactions. SIAM Journal on Applied Mathematics, 80(4):1751–1777, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hsu:2020:HPP

[80] Sze-Bi Hsu, Zhihua Liu, and Pierre Magal. A Holling predator–prey model with handling and searching predators. SIAM Journal on Applied Mathematics, 80(4):1778–1795, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2020:BAG

[81] Qianqian Zhang, Biao Tang, Tianyu Cheng, and Sanyi Tang. Bifurcation analysis of a generalized impulsive Kolmogorov model with applications to pest and disease control. SIAM Journal on Applied Mathematics, 80 (4):1796–1819, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Huang:2020:SDT

[82] Zhe Huang and Chunhua Ou. Speed determinacy of traveling waves to a stream-population model with Allee effect. SIAM Journal on Applied Mathematics, 80(4):1820–1840, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chiarello:2020:MML

[83] Felisia A. Chiarello, Jan Friedrich, Paola Goatin, and Simone Göttlich. Micromacro limit of a nonlocal generalized aw-Rascle type model. SIAM Journal on Applied Mathematics, 80(4): 1841–1861, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parkinson:2020:MOH

[84] Christian Parkinson, David Arnold, Andrea Bertozzi, and Stanley Osher. A model for optimal human navigation with stochastic effects. SIAM Journal on Applied Mathematics, 80(4): 1862–1881, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ferreira:2020:AEI

[85] José A. Ferreira, Paula de Oliveira, and Luís Pinto. Aging effect on iontophoretic transdermal drug delivery. SIAM Journal on Applied Mathematics, 80(4):1882–1907, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mofidi:2020:RPR

[86] Hamid Mofidi and Weishi Liu. Reversal potential and reversal permanent charge with unequal diffusion coefficients via classical Poisson–Nernst–Planck models. SIAM Journal on Applied Mathematics, 80(4):1908–1935, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boros:2020:WRM

[87] Balázs Boros, Gheorghe Craciun, and Polly Y. Yu. Weakly reversible massaction systems with infinitely many positive steady states. SIAM Journal on Applied Mathematics, 80(4): 1936–1946, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ilmavirta:2020:TCT

[88] Joonas Ilmavirta, Olli Koskela, and Jesse Railo. Torus computed tomography. SIAM Journal on Applied Mathematics, 80(4):1947–1976, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Spiridonov:2020:MNA

[89] Alexander O. Spiridonov, Alina Oktyabrskaya, Evgenii M. Karchevskii, and Alexander I. Nosich. Mathematical and numerical analysis of the generalized complex-frequency eigenvalue problem for two-dimensional optical microcavities. SIAM Journal on Applied Mathematics, 80(4):1977–1998, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chen:2020:EAP

[90] Shanshan Chen and Junping Shi. Erratum: Asymptotic Profiles of Basic Reproduction Number for Epidemic Spreading in Heterogeneous Environment. SIAM Journal on Applied Mathematics, 80(4):1999–2000, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [57].

Jiang:2020:EAS

[91] Yu Jiang, Hiroshi Fujiwara, and Gen Nakamura. Erratum for “Approximate Steady State Models for Magnetic Resonance Elastography”. SIAM Journal on Applied Mathematics, 80(4):2001, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [1].

Li:2020:GDH

[92] Chin-Lung Li and Jinn-Liang Liu. Generalized Debye–Hückel equation from Poisson–Bikerman theory. SIAM Journal on Applied Mathematics, 80(5): 2003–2023, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garcia-Cervera:2020:BVF

[93] Carlos J. García-Cervera, Tiziana Giorgi, and Sookyung Joo. Boundary vortex formation in polarizationmodulated orthogonal smectic liquid crystals. SIAM Journal on Applied Mathematics, 80(5):2024–2044, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lin:2020:MTF

[94] Junshan Lin, Stephen P. Shipman, and Hai Zhang. A mathematical theory for Fano resonance in a periodic array of narrow slits. SIAM Journal on Applied Mathematics, 80(5): 2045–2070, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ciesielski:2020:ADB

[95] Danielle Ciesielski, Stephanie McCalla, and Tomas Gedeon. Analysis of dynamics of a biphasic isothermal DNA amplification reaction. SIAM Journal on Applied Mathematics, 80(5): 2071–2097, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zemlyanova:2020:IMP

[96] Anna Y. Zemlyanova. Interaction of multiple plane straight fractures in the presence of the Steigmann–Ogden surface energy. SIAM Journal on Applied Mathematics, 80(5):2098–2119, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nazaikinskii:2020:ESD

[97] Vladimir E. Nazaikinskii, Pavel G. Bedrikovetsky, Liudmila I. Kuzmina, and Yuri V. Osipov. Exact solution for deep bed filtration with finite blocking time. SIAM Journal on Applied Mathematics, 80(5):2120–2143, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gao:2020:HDD

[98] Daozhou Gao. How does dispersal affect the infection size? SIAM Journal on Applied Mathematics, 80 (5):2144–2169, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Reischmann:2020:MMS

[99] L. Reischmann and M. A. Peter. Multiscale modeling and simulation of a Cahn–Larché system with phase separation on the microscale. SIAM Journal on Applied Mathematics, 80 (5):2170–2193, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Boegli:2020:EMM

[100] Sabine Boegli and Christiane Tretter. Eigenvalues of magnetohydrodynamic mean-field dynamo models: Bounds and reliable computation. SIAM Journal on Applied Mathematics, 80(5): 2194–2225, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fikl:2020:JRC

[101] Alexandru Fikl and Daniel J. Bodony. Jump relations of certain hypersingular Stokes kernels on regular surfaces.

SIAM Journal on Applied Mathematics, 80(5):2226–2248, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Natchiar:2020:ACR

[102] S. R. Monisha Natchiar, Richard E. Hewitt, Phillip D. D. Monks, and Peter Morrall. Asymptotics of coupled reaction–diffusion fronts with multiple static and diffusing reactants: Uranium oxidation in water vapor. SIAM Journal on Applied Mathematics, 80 (5):2249–2270, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2020:AFM

[103] Bo Zhang and Haiwen Zhang. An approximate factorization method for inverse acoustic scattering with phaseless total-field data. SIAM Journal on Applied Mathematics, 80(5):2271–2298, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dai:2020:MCH

[104] Shibin Dai, Bo Li, and Toai Luong.

Minimizers for the Cahn–Hilliard energy functional under strong anchoring conditions. SIAM Journal on Applied Mathematics, 80(5):2299–2317, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fan:2020:ECS

[105] Haitao Fan and Chi-Wang Shu. Existence and computation of solutions of a model of traffic involving hysteresis. SIAM Journal on Applied Mathematics, 80(6):2319–2337, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lindstrom:2020:DSM

[106] Torsten Lindström, Yuanji Cheng, and Subhendu Chakraborty. Destabilization, stabilization, and multiple attractors in saturated mixotrophic environments. SIAM Journal on Applied Mathematics, 80(6):2338–2364, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2020:NFC

[107] Luchan Zhang and Yang Xiang. A new formulation of coupling and sliding motions of grain boundaries based on dislocation structure. SIAM Journal on Applied Mathematics, 80(6): 2365–2387, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fan:2020:NHM

[108] Yuwei Fan, Ruo Li, and Lingchao Zheng. A nonlinear hyperbolic model for radiative transfer equation in slab geometry. SIAM Journal on Applied Mathematics, 80(6):2388–2419, ???? 2020. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Roques:2020:AGT

[109] Lionel Roques, Florian Patout, Olivier Bonnefon, and Guillaume Martin. Adaptation in general temporally changing environments. SIAM Journal on Applied Mathematics, 80(6):2420–2447, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kwiecinski:2020:IAI

[110] James A. Kwiecinski, Alain Goriely, and S. Jon Chapman. Interactions of anisotropic inclusions on a fluid membrane. SIAM Journal on Applied Mathematics, 80(6):2448–2471, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jenssen:2020:ABR

[111] Helge Kristian Jenssen and Charis Tsikkou. Amplitude blowup in radial isentropic Euler flow. SIAM Journal on Applied Mathematics, 80 (6):2472–2495, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Jiang:2020:SPN

[112] Tianpeng Jiang, Yang Xiang, and Luchan Zhang. Stochastic Peierls– Nabarro model for dislocations in high entropy alloys. SIAM Journal on Applied Mathematics, 80(6):2496–2517, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Xu:2020:CLA

[113] Jie Xu. Classifying local anisotropy formed by rigid molecules: Symmetries and tensors. SIAM Journal on Applied Mathematics, 80(6):2518–2546, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:2020:SAI

[114] Peijun Li, Jian Zhai, and Yue Zhao.

Stability for the acoustic inverse source problem in inhomogeneous media. SIAM Journal on Applied Mathematics, 80 (6):2547–2559, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Uhlmann:2020:CNN

[115] Gunther Uhlmann and Yiran Wang. Convolutional neural networks in phase space and inverse problems. SIAM Journal on Applied Mathematics, 80 (6):2560–2585, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Evans:2020:TPB

[116] Ryan M. Evans, Arvind Balijepalli, and Anthony J. Kearsley. Transport phenomena in biological field effect transistors. SIAM Journal on Applied Mathematics, 80(6):2586–2607, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wang:2020:PPP

[117] Zhenkun Wang and Hao Wang. Persistence and propagation of a PDE and discrete-time map hybrid animal movement model with habitat shift driven by climate change. SIAM Journal on Applied Mathematics, 80(6):2608–2630, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Amorim:2020:PPI

[118] Paulo Amorim. Predator–prey interactions with hunger structure. SIAM Journal on Applied Mathematics, 80 (6):2631–2656, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barucq:2020:EAA

[119] Hélène Barucq, Florian Faucher, Damien Fournier, Laurent Gizon, and Ha Pham. Efficient and accurate algorithm for the full modal Green’s kernel of the scalar wave equation in helioseismology. SIAM Journal on Applied Mathematics, 80(6):2657–2683, ???? 2020. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lawley:2021:EFI

[120] Sean D. Lawley. The effects of fast inactivation on conditional first passage times of mortal diffusive searchers. SIAM Journal on Applied Mathematics, 81(1):1–24, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:2021:ITH

[121] Guang-Hui Hu, Manmohan Vashisth, and Jiaqing Yang. Inverse timeharmonic electromagnetic scattering from coated polyhedral scatterers with a single far-field pattern. SIAM Journal on Applied Mathematics, 81(1):25–46, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gordon:2021:TFI

[122] Peter V. Gordon, Uday G. Hegde, and Michael C. Hicks. On traveling front of ignition in co-flow laminar reactive jets. SIAM Journal on Applied Mathematics, 81(1):47–59, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Assier:2021:SOQ

[123] Raphael C. Assier and I. David Abrahams. A surprising observation in the quarter-plane diffraction problem. SIAM Journal on Applied Mathematics, 81(1):60–90, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klingenberg:2021:REC

[124] Christian Klingenberg, Ru-Yu Lai, and Qin Li. Reconstruction of the emission coefficient in the nonlinear radiative transfer equation. SIAM Journal on Applied Mathematics, 81(1):91–106, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bessonov:2021:EDS

[125] Nikolai Bessonov, Gennady Bocharov, Andreas Meyerhans, Vladimir Popov, and Vitaly Volpert. Existence and dynamics of strains in a nonlocal reaction– diffusion model of viral evolution. SIAM Journal on Applied Mathematics, 81 (1):107–128, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Philip:2021:MES

[126] Rachel M. Philip, Ian J. Hewitt, and Peter D. Howell. Modeling the effect of surfactant on droplet breakup in a turbulent flow. SIAM Journal on Applied Mathematics, 81(1):129–152, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mazari:2021:FPN

[127] Idriss Mazari and Domènec Ruiz-Balet. A fragmentation phenomenon for a nonenergetic optimal control problem: Optimization of the total population size in logistic diffusive models. SIAM Journal on Applied Mathematics, 81 (1):153–172, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Laurent-Brouty:2021:MTF

[128] Nicolas Laurent-Brouty, Guillaume Costeseque, and Paola Goatin. A macroscopic traffic flow model accounting for bounded acceleration. SIAM Journal on Applied Mathematics, 81(1):173–189, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lee:2021:CPE

[129] Wonjun Lee, Siting Liu, Hamidou Tembine, Wuchen Li, and Stanley Osher. Controlling propagation of epidemics via mean-field control. SIAM Journal on Applied Mathematics, 81 (1):190–207, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Vynnycky:2021:NTR

[130] Michael Vynnycky, Sean McKee, and Lesław Bieniasz. A NonLinear transient reaction–diffusion problem from electroanalytical chemistry. SIAM Journal on Applied Mathematics, 81(1):208–232, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zhang:2021:SDT

[131] Ruming Zhang. Spectrum decomposition of translation operators in periodic waveguide. SIAM Journal on Applied Mathematics, 81(1):233–257, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gottlich:2021:SOT

[132] Simone Göttlich, Michael Herty, Salissou Moutari, and Jennifer Weissen. Second-order traffic flow models on networks. SIAM Journal on Applied Mathematics, 81(1):258–281, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kirsch:2021:EIP

[133] Andreas Kirsch and Andreas Rieder. Erratum: Inverse Problems for Abstract Evolution Equations II: Higher Order Differentiability for Viscoelasticity. SIAM Journal on Applied Mathematics, 81(1):282–283, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [3].

Pedersen:2021:ACE

[134] Geir K. Pedersen. Asymptotic, convergent, and exact truncating series solutions of the linear shallow water equations for channels with power law geometry. SIAM Journal on Applied Mathematics, 81(2):285–303, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gartland:2021:EFI

[135] Eugene C. Gartland, Jr. Electric-field-induced instabilities in nematic liquid crystals. SIAM Journal on Applied Mathematics, 81(2):304–334, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kim:2021:MHD

[136] Yong-Jung Kim and Hyowon Seo. Model for heterogeneous diffusion. SIAM Journal on Applied Mathematics, 81 (2):335–354, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Collin:2021:JSP

[137] Annabelle Collin, Thibaut Kritter, Clair Poignard, and Olivier Saut. Joint stateparameter estimation for tumor growth model. SIAM Journal on Applied Mathematics, 81(2):355–377, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gomez:2021:AAL

[138] Daniel Gomez, Michael J. Ward, and Juncheng Wei. An asymptotic analysis of localized three-dimensional spot patterns for the Gierer–Meinhardt model: Existence, linear stability, and slow dynamics. SIAM Journal on Applied Mathematics, 81(2):378–406, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cosner:2021:EDM

[139] Chris Cosner and Nancy Rodriguez.

The effect of directed movement on the strong Allee effect. SIAM Journal on Applied Mathematics, 81(2):407–433, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Villa:2021:MEP

[140] Chiara Villa, Mark A. Chaplain, and Tommaso Lorenzi. Modeling the emergence of phenotypic heterogeneity in vascularized tumors. SIAM Journal on Applied Mathematics, 81(2):434–453, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cances:2021:GCM

[141] Clément Cancès and David Maltese. A gravity current model with capillary trapping for oil migration in multilayer geological basins. SIAM Journal on Applied Mathematics, 81(2):454–484, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:2021:IRS

[142] Peijun Li and Xu Wang. Inverse random source scattering for the Helmholtz equation with attenuation. SIAM Journal on Applied Mathematics, 81 (2):485–506, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lauretano:2021:RBL

[143] Giulio Lauretano and Renato Spigler. On the relation between the Lifshitz– Slyozov and the Lifshitz–Slyozov– Wagner models for supersaturated solutions. SIAM Journal on Applied Mathematics, 81(2):507–529, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Breit:2021:NMD

[144] Markus Breit and Gillian Queisser. The necessary modeling detail for neuronal signaling: Poisson–Nernst–Planck and cable equation models in one and three dimensions. SIAM Journal on Applied Mathematics, 81(2):530–550, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gonzalez:2021:TSH

[145] Oscar Gonzalez. Theorems on the Stokesian hydrodynamics of a rigid filament in the limit of vanishing radius.

SIAM Journal on Applied Mathematics, 81(2):551–573, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Babb:2021:PLT

[146] T. Babb, G. P. Benham, J. Bows, R. Gonzalez-Farina, K. B. Kiradjiev, W. T. Lee, and S. Tibos. Predicting liftoff time when deep-frying potato dough snacks. SIAM Journal on Applied Mathematics, 81(2):574–590, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kiradjiev:2021:HMR

[147] Kristian B. Kiradjiev, Christopher J.

Breward, Ian Griffiths, and Donald W. Schwendeman. A homogenized model for a reactive filter. SIAM Journal on Applied Mathematics, 81(2):591–619, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Barel:2021:PGA

[148] Ariel Barel, Thomas Dagès, Rotem Manor, and Alfred M. Bruckstein. Probabilistic gathering of agents with simple sensors. SIAM Journal on Applied Mathematics, 81(2):620–640, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ndenda:2021:FOM

[149] Joseph P. Ndenda, John Boscoh H. Njagarah, and Conrad B. Tabi. Fractionalorder model for myxomatosis transmission dynamics: Significance of contact, vector control and culling. SIAM Journal on Applied Mathematics, 81 (2):641–665, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lu:2021:MAW

[150] Wangtao Lu. Mathematical analysis of wave radiation by a step-like surface. SIAM Journal on Applied Mathematics, 81(2):666–693, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Rauch:2021:CGS

[151] Jeffrey Rauch and L. Ridgway Scott. The charge-group summation method for electrostatics of periodic crystals. SIAM Journal on Applied Mathematics, 81(2):694–717, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zheng:2021:MAI

[152] Bo Zheng, Jianshe Yu, and Jia Li. Modeling and analysis of the implementation of the wolbachia incompatible and sterile insect technique for mosquito population suppression. SIAM Journal on Applied Mathematics, 81(2):718–740, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mentus:2021:OMT

[153] Cassidy Mentus and Marcus Roper. Optimal mixing in transport networks: Numerical optimization and analysis. SIAM Journal on Applied Mathematics, 81(3):741–764, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Timms:2021:ARL

[154] Robert Timms, Scott G. Marquis, Valentin Sulzer, Colin P. Please, and S. Jonathan Chapman. Asymptotic reduction of a lithium-ion pouch cell model. SIAM Journal on Applied Mathematics, 81(3):765–788, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cantrell:2021:IFD

[155] Robert Stephen Cantrell, Chris Cosner, and King-Yeung Lam. Ideal free dispersal under general spatial heterogeneity and time periodicity. SIAM Journal on Applied Mathematics, 81 (3):789–813, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Garnier:2021:PCA

[156] Josselin Garnier. Passive communication with ambient noise. SIAM Journal on Applied Mathematics, 81 (3):814–833, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Skene:2021:TLT

[157] David M. Skene and Luke G. Bennetts. A transition-loss theory for waves reflected and transmitted by an overwashed body. SIAM Journal on Applied Mathematics, 81(3):834–852, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Friesen:2021:SWF

[158] Martin Friesen, Hanno Gottschalk, Barbara Rüdiger, and Antoine Tordeux. Spontaneous wave formation in stochastic self-driven particle systems. SIAM Journal on Applied Mathematics, 81 (3):853–870, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nicholls:2021:SLS

[159] David P. Nicholls and Xin Tong. Simulation of localized surface plasmon resonances in two dimensions via impedance–impedance operators. SIAM Journal on Applied Mathematics, 81 (3):871–896, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bevilacqua:2021:MMT

[160] Giulia Bevilacqua, Pasquale Ciarletta, and Alfio Quarteroni. Morphomechanical model of the torsional Clooping in the embryonic heart. SIAM Journal on Applied Mathematics, 81 (3):897–918, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bressloff:2021:DSC

[161] Paul C. Bressloff. Directional searchand-capture model of cytoneme-based morphogenesis. SIAM Journal on Applied Mathematics, 81(3):919–938, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Albuquerque:2021:LSB

[162] Yuri F. Albuquerque, Antoine Laurain, and Irwin Yousept. Level setbased shape optimization approach for sharp-interface reconstructions in timedomain full waveform inversion. SIAM Journal on Applied Mathematics, 81 (3):939–964, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Lewis:2021:EET

[163] Owen L. Lewis and James P. Keener. Enhanced electrodiffusive transport across a mucus layer. SIAM Journal on Applied Mathematics, 81(3):965–981, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cox:2021:SMN

[164] Alexander M. Cox, Emma Horton, Andreas E. Kyprianou, and Denis Villemonais. Stochastic methods for neutron transport equation III: Generational many-to-one and keff. SIAM Journal on Applied Mathematics, 81 (3):982–1001, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Pandian:2021:SDS

[165] Arun Pandian, Jiahe Tony Li, and Snezhana I. Abarzhi. Scale-dependent and self-similar Rayleigh–Taylor and Richtmyer–Meshkov dynamics induced by acceleration varying with length scale. SIAM Journal on Applied Mathematics, 81(3):1002–1019, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ando:2021:SLP

[166] Kazunori Ando, Hyeonbae Kang, Yoshihisa Miyanishi, and Takashi Nakazawa. Surface localization of plasmons in three dimensions and convexity. SIAM Journal on Applied Mathematics, 81(3): 1020–1033, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Chapman:2021:HFT

[167] S. Jonathan Chapman and Zachary M.

Wilmott. Homogenization of flow through periodic networks. SIAM Journal on Applied Mathematics, 81 (3):1034–1051, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Liu:2021:TUB

[168] Jian-Guo Liu, Min Tang, Li Wang, and Zhennan Zhou. Toward understanding the boundary propagation speeds in tumor growth models. SIAM Journal on Applied Mathematics, 81(3): 1052–1076, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yousefnezhad:2021:OCB

[169] Mohsen Yousefnezhad, Chiu-Yen Kao, and Seyyed Abbas Mohammadi. Optimal chemotherapy for brain tumor growth in a reaction-diffusion model. SIAM Journal on Applied Mathematics, 81(3):1077–1097, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Fan:2021:GRP

[170] Chao Fan, Bo Li, and Michael R. White. A generalized Rayleigh–Plesset equation for ions with solvent fluctuations. SIAM Journal on Applied Mathematics, 81(3):1098–1115, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Khademi:2021:AOA

[171] Amin Khademi and Xin Liu. Asymptotically optimal allocation policies for transplant queueing systems. SIAM Journal on Applied Mathematics, 81 (3):1116–1140, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Courtais:2021:SOF

[172] Alexis Courtais, Abderrazak M. Latifi, François Lesage, and Yannick Privat. Shape optimization of fixed-bed reactors in process engineering. SIAM Journal on Applied Mathematics, 81 (3):1141–1165, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kirkland:2021:IVC

[173] Stephen Kirkland, Zhisheng Shuai, P. van den Driessche, and Xueying Wang. Impact of varying community networks on disease invasion. SIAM Journal on Applied Mathematics, 81 (3):1166–1189, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kodsi:2021:BVP

[174] Costy Kodsi and Andrey P. Jivkov.

Boundary value problem on a weighted graph relevant to the static analysis of truss structures. SIAM Journal on Applied Mathematics, 81(3): 1190–1201, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Choi:2021:SPB

[175] Sun-Ho Choi and Hyowon Seo. Synchronization in a power balance system with inertia and nonlinear derivatives. SIAM Journal on Applied Mathematics, 81(3):1202–1225, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ferreira:2021:LSL

[176] J. A. Ferreira, Paula de Oliveira, P. M.

da Silva, and M. Grassi. From lifesaving to life-threatening: a mathematical model to simulate bacterial infections in surgical procedures. SIAM Journal on Applied Mathematics, 81 (3):1226–1247, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Potts:2021:SSS

[177] Jonathan R. Potts and Kevin J. Painter. Stable steady-state solutions of some biological aggregation models. SIAM Journal on Applied Mathematics, 81 (3):1248–1263, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Nouaoura:2021:MAT

[178] Sarra Nouaoura, Nahla Abdellatif, Radhouane Fekih-Salem, and Tewfik Sari. Mathematical analysis of a three-tiered model of anaerobic digestion. SIAM Journal on Applied Mathematics, 81 (3):1264–1286, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Stana:2021:DDC

[179] Remus Stana, Grant Lythe, and Carmen Molina-París. Diffusion in a disk with a circular inclusion. SIAM Journal on Applied Mathematics, 81 (3):1287–1302, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ackley:2021:DKP

[180] Matthew Ackley and Peter Stechlinski. Determining key parameters in riots using lexicographic directional differentiation. SIAM Journal on Applied Mathematics, 81(3):1303–1331, ???? 2021.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Griesmaier:2021:EMI

[181] Roland Griesmaier and Bastian Harrach. Erratum: Monotonicity in inverse medium scattering on unbounded domains. SIAM Journal on Applied Mathematics, 81(3):1332–1337, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [2].

Yang:2021:ANS

[182] Chang Yang and Léon Matar Tine.

Analysis and numerical simulations of a reaction–diffusion model with fixed active bodies. SIAM Journal on Applied Mathematics, 81(4):1339–1360, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

King:2021:MAN

[183] John R. King, Jakub Köry, and Mariya Ptashnyk. Multiscale analysis of nutrient uptake by plant roots with sparse distribution of root hairs: Nonstandard scaling. SIAM Journal on Applied Mathematics, 81(4):1361–1388, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yu:2021:LGS

[184] Yue Yu, Gianmarc Grazioli, Nolan E.

Phillips, and Carter T. Butts. Local graph stability in exponential family random graph models. SIAM Journal on Applied Mathematics, 81 (4):1389–1415, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mahler:2021:ACM

[185] Barbara I. Mahler. Analysis of contagion maps on a class of networks that are spatially embedded in a torus. SIAM Journal on Applied Mathematics, 81(4):1416–1440, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dekens:2021:FPS

[186] Léonard Dekens and Florian Lavigne. Front propagation of a sexual population with evolution of dispersion: a formal analysis. SIAM Journal on Applied Mathematics, 81(4):1441–1460, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kaltenbacher:2021:ILT

[187] Barbara Kaltenbacher and Vanja Nikolić. The inviscid limit of third-order linear and nonlinear acoustic equations. SIAM Journal on Applied Mathematics, 81(4):1461–1482, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bennetts:2021:CRI

[188] L. G. Bennetts and M. H. Meylan. Complex resonant ice shelf vibrations. SIAM Journal on Applied Mathematics, 81(4):1483–1502, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Canic:2021:NGM

[189] Suncica Canić, Yifan Wang, and Martina Bukac. A next-generation mathematical model for drug-eluting stents. SIAM Journal on Applied Mathematics, 81(4):1503–1529, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kow:2021:CNS

[190] Pu-Zhao Kow and Jenn-Nan Wang.

On the characterization of nonradiating sources for the elastic waves in anisotropic inhomogeneous media. SIAM Journal on Applied Mathematics, 81(4):1530–1551, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Koprucki:2021:DHW

[191] Thomas Koprucki, Anieza Maltsi, and Alexander Mielke. On the Darwin– Howie–Whelan equations for the scattering of fast electrons described by the Schrödinger equation. SIAM Journal on Applied Mathematics, 81(4): 1552–1578, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Feng:2021:RDS

[192] Tao Feng, Zhipeng Qiu, and Yun Kang. Recruitment dynamics of social insect colonies. SIAM Journal on Applied Mathematics, 81(4):1579–1599, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dong:2021:PSS

[193] Fang-Di Dong, Jin Shang, William Fagan, and Bingtuan Li. Persistence and spread of solutions in a two-species Lotka–Volterra competitiondiffusion model with a shifting habitat. SIAM Journal on Applied Mathematics, 81(4):1600–1622, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Joshi:2021:ANI

[194] Badal Joshi and Gheorghe Craciun. Autocatalytic networks: an intimate relation between network topology and dynamics. SIAM Journal on Applied Mathematics, 81(4):1623–1644, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ma:2021:MPN

[195] Manman Ma, Zhenli Xu, and Liwei Zhang. Modified Poisson–Nernst– Planck model with Coulomb and hardsphere correlations. SIAM Journal on Applied Mathematics, 81(4):1645–1667, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yaman:2021:IPP

[196] Olha Ivanyshyn Yaman and Frédérique Le Louër. An inverse parameter problem with generalized impedance boundary condition for two-dimensional linear viscoelasticity. SIAM Journal on Applied Mathematics, 81(4):1668–1690, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gorder:2021:LBI

[197] Robert A. Van Gorder, Alissa Kamilova, Rolf G. Birkeland, and Andrew L. Krause. Locating the baking isotherm in a Søderberg electrode: Analysis of a moving thermistor model. SIAM Journal on Applied Mathematics, 81 (4):1691–1716, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Polk:2021:NWD

[198] Sam L. Polk and Bruce M. Boghosian.

The nonuniversality of wealth distribution tails near wealth condensation criticality. SIAM Journal on Applied Mathematics, 81(4):1717–1741, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cassidy:2021:DDD

[199] Tyler Cassidy. Distributed delay differential equation representations of cyclic differential equations. SIAM Journal on Applied Mathematics, 81(4): 1742–1766, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Turzi:2021:DDT

[200] Stefano Turzi. Drift-diffusion transport in a randomly inhomogeneous onedimensional medium. SIAM Journal on Applied Mathematics, 81(4): 1767–1780, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

El-Morshedy:2021:ACD

[201] Hassan A. El-Morshedy and Alfonso Ruiz-Herrera. Asymptotic convergence in delay differential equations arising in epidemiology and physiology. SIAM Journal on Applied Mathematics, 81 (4):1781–1798, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bao:2021:ISP

[202] Gang Bao, Xiang Xu, and Jian Zhai.

Inverse spectral problem for a damped wave operator. SIAM Journal on Applied Mathematics, 81(5):1799–1820, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Robert:2021:SMNa

[203] Philippe Robert and Gaëtan Vignoud. Stochastic models of neural synaptic plasticity. SIAM Journal on Applied Mathematics, 81(5):1821–1846, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Parshad:2021:SNR

[204] Rana D. Parshad, Kwadwo Antwi-Fordjour, and Eric M. Takyi. Some novel results in two species competition. SIAM Journal on Applied Mathematics, 81(5):1847–1869, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Wentz:2021:BCS

[205] Jacqueline M. Wentz and David M. Bortz. Boundedness of a class of spatially discrete reaction–diffusion systems. SIAM Journal on Applied Mathematics, 81(5):1870–1892, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Forien:2021:EMV

[206] Raphaël Forien, Guodong Pang, and Étienne Pardoux. Epidemic models with varying infectivity. SIAM Journal on Applied Mathematics, 81(5): 1893–1930, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Betermin:2021:TFO

[207] Laurent Bétermin. Theta functions and optimal lattices for a grid cells model.

SIAM Journal on Applied Mathematics, 81(5):1931–1953, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Klibanov:2021:LLI

[208] Michael V. Klibanov, Jingzhi Li, and Wenlong Zhang. Linear lavrent’ev integral equation for the numerical solution of a nonlinear coefficient inverse problem. SIAM Journal on Applied Mathematics, 81(5):1954–1978, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Arens:2021:MEC

[209] Tilo Arens, Roland Griesmaier, and Marvin Knöller. Maximizing the electromagnetic chirality of thin dielectric tubes. SIAM Journal on Applied Mathematics, 81(5):1979–2006, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bueler:2021:CLF

[210] Ed Bueler. Conservation laws for freeboundary fluid layers. SIAM Journal on Applied Mathematics, 81(5): 2007–2032, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hansen:2021:ODD

[211] Jakob Hansen and Robert Ghrist. Opinion dynamics on discourse sheaves. SIAM Journal on Applied Mathematics, 81(5):2033–2060, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schreyer:2021:MRM

[212] Lynn Schreyer, Nikos Voulgarakis, Zachary Hilliard, Sergey Lapin, and Loren Cobb. Modeling refugee movement based on a continuum mechanics phase-field approach of porous media. SIAM Journal on Applied Mathematics, 81(5):2061–2082, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bertoglio:2021:RPM

[213] Cristobal Bertoglio, David Nolte, Grigory Panasenko, and Konstantinas Pileckas. Reconstruction of the pressure in the method of asymptotic partial decomposition for the flows in tube structures. SIAM Journal on Applied Mathematics, 81(5):2083–2110, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dong:2021:ACP

[214] Jiu-Gang Dong. Avoiding collisions and pattern formation in flocks. SIAM Journal on Applied Mathematics, 81 (5):2111–2129, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ha:2021:DSA

[215] Seung-Yeal Ha and Hansol Park. A dynamical systems approach for the shape matching of polytopes along rigidbody motions. SIAM Journal on Applied Mathematics, 81(5):2130–2152, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:2021:DFG

[216] Ruo Li, Weiming Li, and Lingchao Zheng. Direct flux gradient approximation to moment closure of kinetic equations. SIAM Journal on Applied Mathematics, 81(5):2153–2179, ???? 2021.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Zabarankin:2021:IPD

[217] Michael Zabarankin and Ying Zhang.

Inverse problem for drop deformation in nonlinear electrohydrodynamics. SIAM Journal on Applied Mathematics, 81 (5):2180–2194, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Schreiber:2021:EQS

[218] Sebastian J. Schreiber, Shuo Huang, Jifa Jiang, and Hao Wang. Extinction and quasi-stationarity for discretetime, endemic SIS and SIR models. SIAM Journal on Applied Mathematics, 81(5):2195–2217, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hodneland:2021:WPD

[219] Erlend Hodneland, Xiaozhe Hu, and Jan M. Nordbotten. Well-posedness and discretization for a class of models for mixed-dimensional problems with high-dimensional gap. SIAM Journal on Applied Mathematics, 81(5): 2218–2245, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Gorb:2021:AEH

[220] Yuliya Gorb and Yuri Kuznetsov.

Asymptotic expansions for high-contrast scalar and vectorial PDEs. SIAM Journal on Applied Mathematics, 81 (5):2246–2264, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Oblakova:2021:RSC

[221] A. Oblakova, A. Al Hanbali, R. J. Boucherie, J. C. W. van Ommeren, and W. H. M. Zijm. Roots, symmetry, and contour integrals in queuing-type systems. SIAM Journal on Applied Mathematics, 81(5):2265–2295, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kawano:2021:DPS

[222] Alexandre Kawano, Antonino Morassi, and Ramón Zaera. Detecting a prey in a spider orb-Web from in-plane vibration. SIAM Journal on Applied Mathematics, 81(6):2297–2322, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Petit:2021:RWD

[223] Julien Petit, Renaud Lambiotte, and Timoteo Carletti. Random walks on dense graphs and graphons. SIAM Journal on Applied Mathematics, 81 (6):2323–2345, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Coclite:2021:GEE

[224] G. M. Coclite, F. Maddalena, G. Puglisi, M. Romano, and G. Saccomandi. The Gardner equation in elastodynamics. SIAM Journal on Applied Mathematics, 81(6):2346–2361, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Robert:2021:SMNb

[225] Philippe Robert and Gaëtan Vignoud. Stochastic models of neural synaptic plasticity: a scaling approach. SIAM Journal on Applied Mathematics, 81 (6):2362–2386, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ji:2021:SRM

[226] Xia Ji and Xiaodong Liu. Source reconstruction with multifrequency sparse scattered fields. SIAM Journal on Applied Mathematics, 81(6):2387–2404, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Tao:2021:DDN

[227] Yiwen Tao, Sue Ann Campbell, and Francis J. Poulin. Dynamics of a diffusive nutrient-phytoplankton-zooplankton model with spatio-temporal delay. SIAM Journal on Applied Mathematics, 81(6):2405–2432, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Levitin:2021:IEL

[228] Michael Levitin, Peter Monk, and Virginia Selgas. Impedance eigenvalues in linear elasticity. SIAM Journal on Applied Mathematics, 81(6):2433–2456, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Alexander:2021:RSE

[229] Amanda M. Alexander and Sean D.

Lawley. Reaction–subdiffusion equations with species-dependent movement. SIAM Journal on Applied Mathematics, 81(6):2457–2479, ???? 2021. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cheng:2021:LME

[230] Jiaqi Cheng, Peter Chesson, and Xiaoying Han. The lottery model for ecological competition in nonstationary environments. SIAM Journal on Applied Mathematics, 81(6):2480–2502, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Li:2021:SIS

[231] Peijun Li, Xiaohua Yao, and Yue Zhao. Stability for an inverse source problem of the biharmonic operator. SIAM Journal on Applied Mathematics, 81 (6):2503–2525, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Crowdy:2021:VMF

[232] Darren G. Crowdy. Viscous Marangoni flow driven by insoluble surfactant and the complex Burgers equation. SIAM Journal on Applied Mathematics, 81 (6):2526–2546, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Dang:2021:HNS

[233] Thuyen Dang, Yuliya Gorb, and Silvia Jiménez Bolaños. Homogenization of nondilute suspension of viscous fluid with magnetic particles. SIAM Journal on Applied Mathematics, 81 (6):2547–2568, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Hu:2021:THA

[234] Guanghui Hu, Wangtao Lu, and Andreas Rathsfeld. Time-harmonic acoustic scattering from locally perturbed periodic curves. SIAM Journal on Applied Mathematics, 81(6):2569–2595, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ledoux:2021:ABE

[235] Jeremy Ledoux, Sebastián Riffo, and Julien Salomon. Analysis of the blade element momentum theory. SIAM Journal on Applied Mathematics, 81 (6):2596–2621, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Isaacson:2021:HRD

[236] Samuel A. Isaacson, Jingwei Ma, and Konstantinos Spiliopoulos. How reaction-diffusion PDEs approximate the large-population limit of stochastic particle models. SIAM Journal on Applied Mathematics, 81(6):2622–2657, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Mercier:2021:VMA

[237] Jean-François Mercier. Variational methods for acoustic radiation in a duct with a shear flow and an absorbing boundary. SIAM Journal on Applied Mathematics, 81(6):2658–2683, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Cohen:2021:NMA

[238] Reed B. Cohen and Jiaxu Li. A novel model and its analysis on the metabolic regulations of glucose, insulin, and glucagon. SIAM Journal on Applied Mathematics, 81(6): 2684–2703, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Bottcher:2021:CQR

[239] Lucas Böttcher and Mason A. Porter. Classical and quantum random-walk centrality measures in multilayer networks. SIAM Journal on Applied Mathematics, 81(6):2704–2724, ???? 2021. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See erratum [342].

Deng:2022:GPE

[240] Youjun Deng, Hongyu Liu, Xianchao Wang, and Wei Wu. On geometrical properties of electromagnetic transmission eigenfunctions and artificial mirage. SIAM Journal on Applied Mathematics, 82(1):1–24, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1413547.

Lunz:2022:ONC

[241] Davin Lunz. Optimizing noisy complex systems liable to failure. SIAM Journal on Applied Mathematics, 82(1):25– 48, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1416126.

Lu:2022:DRT

[242] Jianfeng Lu, Jeremy L. Marzuola, and Alexander B. Watson. Defect resonances of truncated crystal structures. SIAM Journal on Applied Mathematics, 82(1): 49–74, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/21M1415601.

Ambrosi:2022:EMM

[243] D. Ambrosi, L. Deorsola, S. Turzi, and M. Zoppello. Elementary mechanics of the mitral valve. SIAM Journal on Applied Mathematics, 82(1):75–94, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1416655.

Zhou:2022:SME

[244] Peng Zhou and Qihua Huang. A spatiotemporal model for the effects of toxicants on populations in a polluted river. SIAM Journal on Applied Mathematics, 82(1):95–118, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1405629.

Li:2022:MRB

[245] Hongjie Li, Hongyu Liu, and Jun Zou.

Minnaert resonances for bubbles in soft elastic materials. SIAM Journal on Applied Mathematics, 82(1):119–141, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1400572.

Zhigun:2022:NDR

[246] Anna Zhigun and Christina Surulescu.

A novel derivation of rigorous macroscopic limits from a micro-meso description of signal-triggered cell migration in fibrous environments. SIAM Journal on Applied Mathematics, 82(1):142–167, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/20M1365442.

Psaltis:2022:HSW

[247] Steven Psaltis, Robert Timms, Colin Please, and S. Jonathan Chapman. Homogenization of spirally wound highcontrast layered materials. SIAM Journal on Applied Mathematics, 82(1):168– 193, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/20M1377904.

Gamba:2022:RTP

[248] Irene M. Gamba, Qin Li, and Anjali Nair. Reconstructing the thermal phonon transmission coefficient at solid interfaces in the phonon transport equation. SIAM Journal on Applied Mathematics, 82(1):194–220, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/20M1381666.

Huang:2022:SNT

[249] Kuang Huang and Qiang Du. Stability of a nonlocal traffic flow model for connected vehicles. SIAM Journal on Applied Mathematics, 82(1):221–243, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/20M1355732.

Liu:2022:MSM

[250] Xiaodong Liu, Shixu Meng, and Bo Zhang. Modified sampling method with near field measurements. SIAM Journal on Applied Mathematics, 82(1): 244–266, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/21M1432235.

Fogelson:2022:DFB

[251] Aaron L. Fogelson, Anna C. Nelson, Cheryl Zapata-Allegro, and James P. Keener. Development of fibrin branch structure before and after gelation. SIAM Journal on Applied Mathematics, 82(1):267–293, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1401024.

Miroshnichenko:2022:PLM

[252] Taisia Miroshnichenko, Vladimir Gubernov, Sergey Minaev, Vladimir Mislavskii, and Junnosuke Okajima. Piecewise linear model of phytoplankton wave propagation in periodical vortex flow. SIAM Journal on Applied Mathematics, 82(1):294–312, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1405861.

Kincl:2022:TAR

[253] Ondrej Kincl, Michal Pavelka, Frantisek Marsík, and Miroslav Sedlácek. Theoretical analysis of rolling fluid turbines. SIAM Journal on Applied Mathematics, 82(1):313–329, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1392796.

Altmann:2022:LDG

[254] Robert Altmann, Patrick Henning, and Daniel Peterseim. Localization and delocalization of ground states of Bose–Einstein condensates under disorder. SIAM Journal on Applied Mathematics, 82(1):330–358, ????

2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/20M1342434.

Ballif:2022:ASS

[255] Guillaume Ballif, Frédérique Clément, and Romain Yvinec. Averaging of a stochastic slow-fast model for population dynamics: Application to the development of ovarian follicles. SIAM Journal on Applied Mathematics, 82(1):359– 380, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/21M1409615.

Liu:2022:QLT

[256] A. Liu and F. M. G. Magpantay. A quantification of long transient dynamics. SIAM Journal on Applied Mathematics, 82(2):381–407, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/20M1367131.

Acosta:2022:NUI

[257] Sebastian Acosta, Gunther Uhlmann, and Jian Zhai. Nonlinear ultrasound imaging modeled by a Westervelt equation. SIAM Journal on Applied Mathematics, 82(2):408–426, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1431813.

Darde:2022:CAE

[258] J. Dardé, N. Hyvönen, T. Kuutela, and T. Valkonen. Contact adapting electrode model for electrical impedance tomography. SIAM Journal on Applied Mathematics, 82(2):427–449, ????

2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1396125.

Brun:2022:MPS

[259] Mats K. Brun, Elyes Ahmed, Jan M.

Nordbotten, and Nils ChrStenseth. Modeling the process of speciation using a multiscale framework including a posteriori error estimates. SIAM Journal on Applied Mathematics, 82(2):450–475, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1405228.

Marchner:2022:CNA

[260] Philippe Marchner, Xavier Antoine, Christophe Geuzaine, and Hadrien Bériot. Construction and numerical assessment of local absorbing boundary conditions for heterogeneous timeharmonic acoustic problems. SIAM Journal on Applied Mathematics, 82(2): 476–501, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/21M1414929.

Ciallella:2022:LDR

[261] Alessandro Ciallella, Emilio N. M. Cirillo, and Barbara Vantaggi. Localization of defects via residence time measures. SIAM Journal on Applied Mathematics, 82(2):502–525, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/20M1380284.

Buze:2022:SFA

[262] Maciej Buze, Thoms E. Woolley, and Angela Mihai. A stochastic framework for atomistic fracture. SIAM Journal on Applied Mathematics, 82(2):526–548, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1416436.

Benes:2022:QNA

[263] Michal Benes, Miroslav Kolár, and Daniel Sevcovic. Qualitative and numerical aspects of a motion of a family of interacting curves in space. SIAM Journal on Applied Mathematics, 82(2):549–575, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1417181.

Liu:2022:NMP

[264] Shuang Liu, Yuan Lou, and Pengfei Song. A new monotonicity for principal eigenvalues with applications to timeperiodic patch models. SIAM Journal on Applied Mathematics, 82(2):576–601, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/20M1320973.

Lai:2022:ITD

[265] Ru-Yu Lai, Kui Ren, and Ting Zhou. Inverse transport and diffusion problems in photoacoustic imaging with nonlinear absorption. SIAM Journal on Applied Mathematics, 82(2):602–624, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1436178.

Kirk:2022:MEH

[266] Toby L. Kirk, Jack Evans, Colin P.

Please, and S. Jonathan Chapman. Modeling electrode heterogeneity in lithium-ion batteries: Unimodal and bimodal particle-size distributions. SIAM Journal on Applied Mathematics, 82(2): 625–653, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/20M1344305.

Chen:2022:TNP

[267] Shanshan Chen, Junping Shi, Zhisheng Shuai, and Yixiang Wu. Two novel proofs of spectral monotonicity of perturbed essentially nonnegative matrices with applications in population dynamics. SIAM Journal on Applied Mathematics, 82(2):654–676, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/20M1345220.

Vynnycky:2022:TNO

[268] Michael Vynnycky and Sean McKee. On a transient nonlinear one-dimensional reaction-diffusion equation with a pointsource initial condition. SIAM Journal on Applied Mathematics, 82(2):677–693, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M140701X.

Dalby:2022:ODF

[269] James Dalby, Patrick E. Farrell, Apala Majumdar, and Jingmin Xia. Onedimensional ferronematics in a channel: Order reconstruction, bifurcations, and multistability. SIAM Journal on Applied Mathematics, 82(2):694–719, ????

2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1400171.

Meng:2022:EMT

[270] Qingle Meng, Zhengjian Bai, Huaian Diao, and Hongyu Liu. Effective medium theory for embedded obstacles in elasticity with applications to inverse problems. SIAM Journal on Applied Mathematics, 82(2):720–749, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1431369.

Fox:2022:HQM

[271] Rodney O. Fox and Frédérique Laurent. Hyperbolic quadrature method of moments for the one-dimensional kinetic equation. SIAM Journal on Applied Mathematics, 82(2):750–771, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1406143.

Tang:2022:AAS

[272] Tao Tang, Boyi Wang, and Jiang Yang. Asymptotic analysis on the sharp interface limit of the time-fractional Cahn– Hilliard equation. SIAM Journal on Applied Mathematics, 82(3):773–792, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1427863.

Park:2022:CGS

[273] Youngmin Park, Prashant Singh, and Thomas G. Fai. Coarse-grained stochastic model of myosin-driven vesicles into dendritic spines. SIAM Journal on Applied Mathematics, 82(3):793–820, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1434180.

Nie:2022:EDD

[274] Hua Nie, Yao Shi, and Jianhua Wu. The effect of diffusion on the dynamics of a predator–prey chemostat model. SIAM Journal on Applied Mathematics, 82(3): 821–848, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/21M1432090.

Negron-Marrero:2022:IEC

[275] Pablo V. Negrón-Marrero and Jeyabal Sivaloganathan. Infinite energy cavitating solutions: a variational approach. SIAM Journal on Applied Mathematics, 82(3):849–871, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1427711.

Nethercote:2022:DAW

[276] Matthew A. Nethercote, Anastasia V.

Kisil, and Raphael C. Assier. Diffraction of acoustic waves by a wedge of point scatterers. SIAM Journal on Applied Mathematics, 82(3):872–898, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1438608.

Gao:2022:EAI

[277] Daozhou Gao, Justin M. W. Munganga, P. van den Driessche, and Lei Zhang. Effects of asymptomatic infections on the spatial spread of infectious diseases. SIAM Journal on Applied Mathematics, 82(3):899–923, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/21M1398434.

Dondl:2022:EMS

[278] Patrick Dondl, Patrina S. P. Poh, and Marius Zeinhofer. An efficient model for scaffold mediated bone regeneration. SIAM Journal on Applied Mathematics, 82(3):924–949, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1401887.

Gu:2022:GRM

[279] Binan Gu, Lou Kondic, and Linda J. Cummings. A graphical representation of membrane filtration. SIAM Journal on Applied Mathematics, 82(3):950–975, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1424743.

Su:2022:GSC

[280] Wei Su and Xiang Zhang. Global stability and Canard explosions of the predator–prey model with the sigmoid functional response. SIAM Journal on Applied Mathematics, 82(3):976–1000, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1437755.

Rastegin:2022:ESE

[281] Alexey Rastegin. Estimating the Shannon entropy and (un)certainty relations for design-structured POVMs. SIAM Journal on Applied Mathematics, 82(3):1001–1019, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1408105.

Cao:2022:BJI

[282] Yining Cao and Xiaoming Wang. The Beavers–Joseph interface boundary condition is well approximated by the Beavers–Joseph–Saffman–Jones interface boundary condition. SIAM Journal on Applied Mathematics, 82(3):1020– 1044, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/21M1462386.

Jin:2022:ROD

[283] Bangti Jin and Yavar Kian. Recovery of the order of derivation for fractional diffusion equations in an unknown medium. SIAM Journal on Applied Mathematics, 82(3):1045–1067, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1398264.

Chao:2022:PRE

[284] Ying Chao and Molei Tao. Parametric resonance for enhancing the rate of metastable transition. SIAM Journal on Applied Mathematics, 82(3):1068–1090, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M144966X.

Werner:2022:DIM

[285] Philipp Werner, Martin Burger, Florian Frank, and Harald Garcke. A diffuse interface model for cell blebbing including membrane–cortex coupling with linker dynamics. SIAM Journal on Applied Mathematics, 82(3):1091–1112, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1433642.

Li:2022:ISP

[286] Jianliang Li, Peijun Li, and Xu Wang. Inverse source problems for the stochastic wave equations: Far-field patterns. SIAM Journal on Applied Mathematics, 82(4):1113–1134, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1467663.

Grabovsky:2022:RSF

[287] Yury Grabovsky. Reconstructing Stieltjes functions from their approximate values: a search for a needle in a haystack. SIAM Journal on Applied Mathematics, 82(4):1135–1166, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1392279.

Wittwer:2022:MPM

[288] Nicolás Alejandro Barnafi Wittwer, Simone Di Gregorio, Luca Dede’, Paolo Zunino, Christian Vergara, and Alfio Quarteroni. A multiscale poromechanics model integrating myocardial perfusion and the epicardial coronary vessels. SIAM Journal on Applied Mathematics, 82(4):1167–1193, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1424482.

Niethammer:2022:OBD

[289] B. Niethammer, R. L. Pego, A. Schlichting, and J. J. L. Velázquez. Oscillations in a Becker–Döring model with injection and depletion. SIAM Journal on Applied Mathematics, 82(4):1194–1219, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/20M1398664.

Schenk:2022:EUN

[290] Christina Schenk and Volker H. Schulz.

Existence, uniqueness, and numerical modeling of wine fermentation based on integro-differential equations. SIAM Journal on Applied Mathematics, 82(4):1220–1245, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/20M1362309.

Dages:2022:DSP

[291] Thomas Dagès and Alfred M. Bruckstein. Doubly stochastic pairwise interactions for agreement and alignment. SIAM Journal on Applied Mathematics, 82(4):1246–1266, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1394680.

Jiang:2022:TGB

[292] Kai Jiang, Wei Si, and Jie Xu. Tilt grain boundaries of hexagonal structures: a spectral viewpoint. SIAM Journal on Applied Mathematics, 82(4):1267–1286, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1463288.

Dhia:2022:HSM

[293] Anne-Sophie Bonnet-Ben Dhia, Simon N. Chandler-Wilde, and Sonia Fliss. On the half-space matching method for real wavenumber. SIAM Journal on Applied Mathematics, 82(4):1287–1311, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/21M1459216.

Willis:2022:HSI

[294] Nathan Willis, Chee Han Tan, Christel Hohenegger, and Braxton Osting. High spots for the ice-fishing problem with surface tension. SIAM Journal on Applied Mathematics, 82(4):1312–1335, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1458879.

Liang:2022:TEC

[295] Nan Liang, Maoxing Liu, Yongzheng Sun, Rui Xiao, and Lingzhi Zhao. Time and energy costs for synchronization of Kuramoto-oscillator networks with or without noise perturbation. SIAM Journal on Applied Mathematics, 82(4):1336–1355, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/21M1457928.

Li:2022:NER

[296] Conway Li, Neville Fowkes, Brendan Florio, and Miccal Matthews.

Near-exact radiating fins via boundary tracing. SIAM Journal on Applied Mathematics, 82(4):1356–1368, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1476228.

Aguayo:2022:AOI

[297] Jorge Aguayo and Hugo Carrillo Lincopi. Analysis of obstacles immersed in viscous fluids using Brinkman’s law for steady Stokes and Navier– Stokes equations. SIAM Journal on Applied Mathematics, 82(4):1369–1386, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/20M138569X.

Schnitzer:2022:APP

[298] Ory Schnitzer and Richard Porter.

Acoustics of a partially partitioned narrow slit connected to a half-plane: Case study for exponential quasi-bound states in the continuum and their resonant excitation. SIAM Journal on Applied Mathematics, 82(4):1387–1410, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1470426.

Li:2022:BTL

[299] Conway Li, Neville Fowkes, and Miccal Matthews. Boundary tracing for Laplace’s equation with conformal mapping. SIAM Journal on Applied Mathematics, 82(4):1411–1422, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1476241.

Ancona:2022:SSA

[300] Fabio Ancona, Alberto Bressan, and Maria Teresa Chiri. Soil searching by an artificial root. SIAM Journal on Applied Mathematics, 82(4):1423–1445, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/20M1338678.

Hiptmair:2022:SQR

[301] Ralf Hiptmair, Andrea Moiola, and Euan A. Spence. Spurious quasiresonances in boundary integral equations for the Helmholtz transmission problem. SIAM Journal on Applied Mathematics, 82(4):1446–1469, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/21M1447052.

Kraitzman:2022:SMB

[302] Noa Kraitzman, Keith Promislow, and Brian Wetton. Slow migration of brine inclusions in first-year sea ice. SIAM Journal on Applied Mathematics, 82(4):1470–1494, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/21M1440244.

Kunz:2022:DRA

[303] Valentin D. Kunz and Raphael Assier. Diffraction by a right-angled no-contrast penetrable wedge revisited: a double Wiener–Hopf approach. SIAM Journal on Applied Mathematics, 82(4):1495– 1519, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/21M1461861.

Davies:2022:CME

[304] Cameron Davies, Tongseok Lim, and Robert J. McCann. Classifying minimum energy states for interacting particles: Spherical shells. SIAM Journal on Applied Mathematics, 82(4):1520–1536, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1455309.

Goldsztein:2022:WPB

[305] Guillermo H. Goldsztein. Wobbling of pedestrian bridges. SIAM Journal on Applied Mathematics, 82(4):1537–1557, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1463148.

Hua:2022:ERO

[306] Yifan Hua, Kevin Miller, Andrea L.

Bertozzi, Chen Qian, and Bao Wang. Efficient and reliable overlay networks for decentralized federated learning. SIAM Journal on Applied Mathematics, 82(4):1558–1586, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1465081.

Sun:2022:DDF

[307] Gui-Quan Sun, Hong-Tao Zhang, Li-Li Chang, Zhen Jin, Hao Wang, and Shigui Ruan. On the dynamics of a diffusive foot-and-mouth disease model with nonlocal infections. SIAM Journal on Applied Mathematics, 82(4):1587–1610, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1412992.

Ammari:2022:VTB

[308] Habib Ammari, Francesco Fiorani, and Erik Orvehed Hiltunen. On the validity of the tight-binding method for describing systems of subwavelength resonators. SIAM Journal on Applied Mathematics, 82(4):1611–1634, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1449804.

Bruna:2022:PSS

[309] Maria Bruna, Martin Burger, Antonio Esposito, and Simon M. Schulz. Phase separation in systems of interacting active Brownian particles. SIAM Journal on Applied Mathematics, 82(4):1635– 1660, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/21M1452524.

Fayolle:2022:SSS

[310] Guy Fayolle, Jean-Marc Lasgouttes, and Carlos Flores. Stability and string stability of car-following models with reaction-time delay. SIAM Journal on Applied Mathematics, 82(5):1661–1679, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1443650.

Wang:2022:SSR

[311] Hao Wang, Kai Wang, and Yong-Jung Kim. Spatial segregation in reaction-diffusion epidemic models. SIAM Journal on Applied Mathematics, 82(5):1680–1709, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1485814.

Zhang:2022:MBF

[312] Wenzhong Zhang, Bo Wang, and Wei Cai. A matrix basis formulation for the dyadic Green’s functions of Maxwell’s equations in layered media. SIAM Journal on Applied Mathematics, 82(5):1710–1732, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1429102.

Iwasaki:2022:SWR

[313] Satoru Iwasaki, Shuichi Jimbo, and Yoshihisa Morita. Standing waves of reaction-diffusion equations on an unbounded graph with two vertices. SIAM Journal on Applied Mathematics, 82(5):1733–1763, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1454572.

Chang:2022:SOC

[314] Lili Chang, Wei Gong, Zhen Jin, and Gui-Quan Sun. Sparse optimal control of pattern formations for an SIR reaction-diffusion epidemic model. SIAM Journal on Applied Mathematics, 82(5):1764–1790, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1472127.

Jahedi:2022:RSA

[315] Sana Jahedi, Timothy Sauer, and James A. Yorke. Robustness of solutions of almost every system of equations. SIAM Journal on Applied Mathematics, 82(5):1791–1807, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1483529.

Shi:2022:NLC

[316] Baoming Shi, Yucen Han, and Lei Zhang. Nematic liquid crystals in a rectangular confinement: Solution landscape, and bifurcation. SIAM Journal on Applied Mathematics, 82(5):1808– 1828, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/21M1447404.

Feng:2022:GCP

[317] Lirui Feng, Yi Wang, and Jianhong Wu. Generic convergence to periodic orbits for monotone flows with respect to 2-cones and applications to SEIRS models. SIAM Journal on Applied Mathematics, 82(5):1829–1850, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1466566.

Cang:2022:SOT

[318] Zixuan Cang, Qing Nie, and Yanxiang Zhao. Supervised optimal transport. SIAM Journal on Applied Mathematics, 82(5):1851–1877, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1469171.

Wang:2022:GDF

[319] Jiafu Wang, Shan Zhang, and Lin Wang. Global dynamics of a Filippov SIQR model with delayed relay control. SIAM Journal on Applied Mathematics, 82(6):1879–1902, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1462830.

Qu:2022:MSW

[320] Zhuolin Qu, Tong Wu, and James M.

Hyman. Modeling spatial waves of wolbachia invasion for controlling mosquitoborne diseases. SIAM Journal on Applied Mathematics, 82(6):1903–1929, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1440384.

Klimm:2022:MMG

[321] Florian Klimm, Nick S. Jones, and Michael T. Schaub. Modularity maximization for graphons. SIAM Journal on Applied Mathematics, 82(6):1930–1952, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1492003.

Lippitt:2022:SIM

[322] William Lippitt, Sunder Sethuraman, and Xueying Tang. Stationarity and inference in multistate promoter models of stochastic gene expression via stickbreaking measures. SIAM Journal on Applied Mathematics, 82(6):1953–1986, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1440876.

Higham:2022:MFA

[323] Desmond John Higham and Henry-Louis de Kergorlay. Mean field analysis of hypergraph contagion models. SIAM Journal on Applied Mathematics, 82(6):1987–2007, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1440219.

Popovic:2022:LDA

[324] Lea Popovic and Giovanni Zoroddu. Large deviations for additive functionals of reflected jump-diffusions. SIAM Journal on Applied Mathematics, 82(6):2008–2035, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1446162.

Sahneh:2022:EEP

[325] Faryad D. Sahneh, William Fries, Joseph C. Watkins, and Joceline Lega. Epidemics from the eye of the pathogen. SIAM Journal on Applied Mathematics, 82(6):2036–2056, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1450719.

Ledger:2022:MOC

[326] P. D. Ledger and W. R. B. Lionheart. Minimal object characterizations using harmonic generalized polarizability tensors and symmetry groups. SIAM Journal on Applied Mathematics, 82(6):2057–2079, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1471547.

Matin:2022:NCD

[327] Hossein Nick Zinat Matin and Richard B.

Sowers. Near-collision dynamics in a noisy car-following model. SIAM Journal on Applied Mathematics, 82 (6):2080–2110, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://

epubs.siam.org/doi/10.1137/21M1454055.

Zhang:2022:CIM

[328] Yaming Zhang, Thomas Stiehl, and Min Tang. A continuous integral model for white blood cell production. SIAM Journal on Applied Mathematics, 82(6):2111–2130, ???? 2022. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M147668X.

Demeio:2023:WPS

[329] Lucio Demeio and Stefano Lenci. Wave propagation on a string resting on a general nonlinear substrate. SIAM Journal on Applied Mathematics, 83(1):1–24, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1503488.

Borcea:2023:PWP

[330] Liliana Borcea, Josselin Garnier, and Knut Sølna. Paraxial wave propagation in random media with long-range correlations. SIAM Journal on Applied Mathematics, 83(1):25–51, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M149524X.

Cheng:2023:EFO

[331] Xinyu Cheng, Zhaohui Fu, and Brian Wetton. Equivalent formulations of the oxygen depletion problem, other implicit moving boundary value problems, and implications for numerical approximation. SIAM Journal on Applied Mathematics, 83(1):52–78, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1419167.

Pelinovsky:2023:KIA

[332] Dmitry Pelinovsky and Guido Schneider. KP-II approximation for a scalar Fermi–Pasta–Ulam system on a 2D square lattice. SIAM Journal on Applied Mathematics, 83(1):79–98, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1509369.

Fang:2023:FRD

[333] Chengran Fang, Demian Wassermann, and Jing-Rebecca Li. Fourier representation of the diffusion MRI signal using layer potentials. SIAM Journal on Applied Mathematics, 83(1):99–121, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1439572.

Klimasara:2023:MSS

[334] Paweł Klimasara and Marta Tyran-Kamińska. A model of seasonal savanna dynamics. SIAM Journal on Applied Mathematics, 83(1):122–143, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1461393.

Lunz:2023:MOC

[335] Davin Lunz and J. Frédéric Bonnans. Modeling and optimal control of a two-species bioproducing microbial consortium. SIAM Journal on Applied Mathematics, 83(1):144–171, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1476113.

Pichor:2023:APG

[336] Katarzyna Pichór and Ryszard Rudnicki. Asymptotic properties of a general model of immune status. SIAM Journal on Applied Mathematics, 83(1):172–193, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1466906.

Putra:2023:FPA

[337] Prama Putra, Hadrien Oliveri, Travis Thompson, and Alain Goriely. Front propagation and arrival times in networks with application to neurodegenerative diseases. SIAM Journal on Applied Mathematics, 83(1):194–224, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1467547.

Schmeller:2023:GFC

[338] Leonie Schmeller and Dirk Peschka.

Gradient flows for coupling order parameters and mechanics. SIAM Journal on Applied Mathematics, 83(1):225–253, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M148478X.

Wei:2023:ENE

[339] Chaozhen Wei and Min Wu. An Eulerian nonlinear elastic model for compressible and fluidic tissue with radially symmetric growth. SIAM Journal on Applied Mathematics, 83(1):254–275, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1463653.

Fan:2023:DMT

[340] Guihong Fan, Juan Li, Jacques Bélair, and Huaiping Zhu. Delayed model for the transmission and control of COVID-19 with fangcang shelter hospitals. SIAM Journal on Applied Mathematics, 83(1):276–301, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M146154X.

Xu:2023:UID

[341] Xiaoxu Xu, Guanghui Hu, Bo Zhang, and Haiwen Zhang. Uniqueness in inverse diffraction grating problems with infinitely many plane waves at a fixed frequency. SIAM Journal on Applied Mathematics, 83(1):302–326, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1480963.

Bottcher:2023:ECQ

[342] Lucas Böttcher and Mason A. Porter. Erratum: Classical and quantum random-walk centrality measures in multilayer networks. SIAM Journal on Applied Mathematics, 83(1):327–328, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1522139.  See

[239].

Bansal:2023:SMN

[343] Diksha Bansal, Dipa Ghosh, and Sarthok Sircar. Selection mechanism in non-Newtonian Saffman–Taylor fingers. SIAM Journal on Applied Mathematics, 83(2):329–353, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1485838.

Britton:2023:OIS

[344] Tom Britton and Lasse Leskelä. Optimal intervention strategies for minimizing total incidence during an epidemic. SIAM Journal on Applied Mathematics, 83(2):354–373, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1504433.

Qiao:2023:MAN

[345] Zhonghua Qiao, Zhenli Xu, Qian Yin, and Shenggao Zhou. A Maxwell– Ampère Nernst–Planck framework for modeling charge dynamics. SIAM Journal on Applied Mathematics, 83(2):374– 393, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/22M1477891.

Meng:2023:SMM

[346] Shixu Meng. Single mode multifrequency factorization method for the inverse source problem in acoustic waveguides. SIAM Journal on Applied Mathematics, 83(2):394–417, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M1462891.

Li:2023:ISP

[347] Wei Li, John C. Schotland, Yang Yang, and Yimin Zhong. Inverse source problem for acoustically-modulated electromagnetic waves. SIAM Journal on Applied Mathematics, 83(2):418–435, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1480690.

Dubovskaya:2023:AMF

[348] Alina Dubovskaya, Susan C. Fennell, Kevin Burke, James P. Gleeson, and Doireann O’Kiely. Analysis of meanfield approximation for Deffuant opinion dynamics on networks. SIAM Journal on Applied Mathematics, 83(2):436–459, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1499765.

Awasthi:2023:EAT

[349] Piyush Awasthi, Mahendhar Kumar, and Yana Nec. Effects of anisotropy in tridimensional diffusion: Flow patterns and transport efficiency. SIAM Journal on Applied Mathematics, 83(2):460–483, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1476642.

Cui:2023:WHF

[350] Jianbo Cui, Shu Liu, and Haomin Zhou.

Wasserstein Hamiltonian flow with common noise on graph. SIAM Journal on Applied Mathematics, 83(2):484–509, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1490697.

Zhang:2023:SEA

[351] Hai-E Zhang, Gen-Qi Xu, and Zhong-Jie Han. Stability and eigenvalue asymptotics of multi-dimensional fully magnetic effected piezoelectric system with friction-type infinite memory. SIAM Journal on Applied Mathematics, 83(2): 510–529, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/22M1486790.

Zhang:2023:IBP

[352] Jimin Zhang, Pingping Cong, and Meng Fan. Interactions between pelagic and benthic producers: Asymmetric competition for light and nutrients. SIAM Journal on Applied Mathematics, 83(2): 530–552, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/22M1491290.

Valovik:2023:MEN

[353] Dmitry V. Valovik. Maxwell’s equations with nonlinear inhomogeneous constitutive relation: Guided waves in a film filled with inhomogeneous Kerr medium. SIAM Journal on Applied Mathematics, 83(2):553–575, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1489587.

Gai:2023:RMC

[354] Chunyi Gai and Theodore Kolokolnikov. Resource-mediated competition between two plant species with different rates of water intake. SIAM Journal on Applied Mathematics, 83(2):576–602, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/21M144623X.

Zemlyanova:2023:APP

[355] Anna Y. Zemlyanova and Lauren M. White. An axisymmetric problem for a patch loaded material surface attached to the boundary of an elastic semi-space. SIAM Journal on Applied Mathematics, 83(2):603–624, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1523753.

Feppon:2023:SRO

[356] Florian Feppon, Zijian Cheng, and Habib Ammari. Subwavelength resonances in one-dimensional high-contrast acoustic media. SIAM Journal on Applied Mathematics, 83(2):625–665, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1503841.

Li:2023:NPM

[357] Buyang Li, Yonglin Li, and Weiying Zheng. A new perfectly matched layer method for the Helmholtz equation in nonconvex domains. SIAM Journal on Applied Mathematics, 83(2):666–694, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1482524.

Katsevich:2023:RDR

[358] Alexander Katsevich. Resolution of 2 dimensional reconstruction of functions with nonsmooth edges from discrete Radon transform data. SIAM Journal on Applied Mathematics, 83(2):695–724, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1466712.

Hanevy:2023:AAF

[359] Niall Hanevy and Doireann O’Kiely.

Asymptotic analysis for fiber drawing processes. SIAM Journal on Applied Mathematics, 83(2):725–747, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1470074.

Banaji:2023:SRP

[360] Murad Banaji. Splitting reactions preserves nondegenerate behaviors in chemical reaction networks. SIAM Journal on Applied Mathematics, 83(2):748–769, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1478392.

Mascali:2023:SET

[361] Giovanni Mascali. Some electric, thermal, and thermoelectric properties of suspended monolayer graphene. SIAM Journal on Applied Mathematics, 83(2): 770–790, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/22M147548X.

Niu:2023:WIH

[362] Yuanling Niu, Hao Kang, Pingyang Wang, Chenchen Guo, Fan Nie, Hongbin Ji, Jiarui Wu, and Luonan Chen. Why increase of heterogeneity signals predeterioration during tumor progression: a unified mathematical model. SIAM Journal on Applied Mathematics, 83(2): 791–815, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/22M1497729.

Glendinning:2023:NFD

[363] Paul A. Glendinning and David J. W. Simpson. Normal forms, differentiable conjugacies, and elementary bifurcations of maps. SIAM Journal on Applied Mathematics, 83(2):816–836, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1503701.

Montefusco:2023:RHL

[364] Alberto Montefusco, Christof Schütte, and Stefanie Winkelmann. A route to the hydrodynamic limit of a reactiondiffusion master equation using gradient structures. SIAM Journal on Applied Mathematics, 83(2):837–861, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/22M1488831.

Bressloff:2023:ATD

[365] Paul C. Bressloff. Accumulation time of diffusion in a 3D singularly perturbed domain. SIAM Journal on Applied Mathematics, 83(2):862–881, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1475843.

Vynnycky:2023:AMG

[366] Michael Vynnycky, Bharath V. Rangavittal, and Björn Glaser. An asymptotic model for gas-solid flow in a countercurrent moving bed reactor. SIAM Journal on Applied Mathematics, 83(2): 882–908, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/22M1496293.

Hayat:2023:DTJ

[367] Amaury Hayat, Benedetto Piccoli, and Sydney Truong. Dissipation of traffic jams using a single autonomous vehicle on a ring road. SIAM Journal on Applied Mathematics, 83(3):909–937, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1414346.

Doak:2023:NSW

[368] Alex Doak, Guido Baardink, Paul A.

Milewski, and Anton Souslov. Nonlinear shallow-water waves with vertical odd viscosity. SIAM Journal on Applied Mathematics, 83(3):938–965, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M149082X.

Miyoshi:2023:GSI

[369] Hiroyuki Miyoshi and Darren G.

Crowdy. Generalized Schwarz integral formulas for multiply connected domains. SIAM Journal on Applied Mathematics, 83(3):966–984, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1506419.

Dosmagulova:2023:OCP

[370] Karlygash Dosmagulova, Vladimir Mityushev, and Zhanat Zhunussova. On the optimal conductivity of packed two-dimensional dispersed composites. SIAM Journal on Applied Mathematics, 83(3):985–999, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1527660.

Worsfold:2023:DFS

[371] Jeremy R. Worsfold, Tim Rogers, and Paul Milewski. Density fluctuations in stochastic kinematic flows. SIAM Journal on Applied Mathematics, 83(3):1000–1024, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1494166.

Hong:2023:CTF

[372] Hyukpyo Hong, Bryan S. Hernandez, Jinsu Kim, and Jae Kyoung Kim. Computational translation framework identifies biochemical reaction networks with special topologies and their longterm dynamics. SIAM Journal on Applied Mathematics, 83(3):1025–1048, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M150469X.

Wang:2023:LMS

[373] Sára Wang, Mirza Karamehmedović, and Faouzi Triki. Localization of moving sources: Uniqueness, stability, and Bayesian inference. SIAM Journal on Applied Mathematics, 83(3):1049–1073, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1500691.

Mozumder:2023:TWC

[374] Ali Mozumder, Artur Safin, Susan E.

Minkoff, and John Zweck. A twoway coupled model of visco-thermo-acoustic effects in photoacoustic trace gas sensors. SIAM Journal on Applied Mathematics, 83(3):1074–1097, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1447659.

Qu:2023:MIM

[375] Zhuolin Qu, Denis Patterson, Lauren M. Childs, Christina J. Edholm, Joan Ponce, Olivia Prosper, and Lihong Zhao. Modeling immunity to malaria with an age-structured PDE framework. SIAM Journal on Applied Mathematics, 83(3):1098–1125, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1464427.

Adriazola:2023:ADE

[376] Jimmie Adriazola and Roy H. Goodman. Apodizer design to efficiently couple light into a fiber Bragg grating. SIAM Journal on Applied Mathematics, 83(3):1126–1145, ???? 2023.

CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1507759.

Celora:2023:DCP

[377] Giulia L. Celora, Matthew G. Hennessy, Andreas Muench, Barbara Wagner, and Sarah Waters. The dynamics of a collapsing polyelectrolyte gel. SIAM Journal on Applied Mathematics, 83(3):1146–1171, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/21M1419726.

Girardin:2023:AEC

[378] Léo Girardin and Baptiste Maucourt. Agro-ecological control of a pesthost system: Preventing spreading. SIAM Journal on Applied Mathematics, 83(3):1172–1195, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1508066.

Gao:2023:IPP

[379] Yixian Gao, Hongyu Liu, and Yang Liu. On an inverse problem for the plate equation with passive measurement. SIAM Journal on Applied Mathematics, 83(3):1196–1214, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1499881.

MacLaurin:2023:PRW

[380] J. MacLaurin. Phase reduction of waves, patterns, and oscillations subject to spatially extended noise. SIAM Journal on Applied Mathematics, 83(3):1215–1244, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1451221.

Zheng:2023:APT

[381] Zhong Zheng. Accumulation of particles towards the envelope of a fluid-infused patch under evaporation. SIAM Journal on Applied Mathematics, 83(3):1245– 1271, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/22M1472784.

Huang:2023:OSC

[382] Wenlin Huang, Jin Liang, and Yuchao Dong. Optimal stochastic control problem for a carbon emission reduction process. SIAM Journal on Applied Mathematics, 83(3):1272–1295, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1496840.

Liu:2023:TVF

[383] X. Liu, J. Song, F. Pourahmadian, and H. Haddar. Time- versus frequencydomain inverse elastic scattering: Theory and experiment. SIAM Journal on Applied Mathematics, 83(3):1296–1314, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1522437.

Bailo:2023:BPF

[384] Rafael Bailo, José A. Carrillo, and Jingwei Hu. Bound-preserving finite-volume schemes for systems of continuity equations with saturation. SIAM Journal on Applied Mathematics, 83(3):1315–1339, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1488703.

Chen:2023:SRP

[385] Junren Chen and Michael K. Ng. Signal reconstruction from phase-only measurements: Uniqueness condition, minimal measurement number and beyond. SIAM Journal on Applied Mathematics, 83(4):1341–1365, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1432077.

Autry:2023:MFR

[386] Eric Autry, Daniel Carter, Gregory J. Herschlag, Zach Hunter, and Jonathan C. Mattingly. Metropolized forest recombination for Monte Carlo sampling of graph partitions. SIAM Journal on Applied Mathematics, 83 (4):1366–1391, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://

epubs.siam.org/doi/10.1137/21M1418010.

Garnier:2023:PDM

[387] Josselin Garnier, Ziyu Lu, and Laurent Mertz. A piecewise deterministic Markov process approach modeling a dry friction problem with noise. SIAM Journal on Applied Mathematics, 83(4):1392–1421, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1480847.

Herrmann:2023:IHP

[388] Michael Herrmann and Barbara Niethammer. Instability of hysteretic phase interfaces in a mean-field model with inhomogeneities. SIAM Journal on Applied Mathematics, 83(4):1422–1443, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M153197X.

Chemnitz:2023:SSS

[389] Robin Chemnitz. Sensitivity of steady states in networks with application to Markov chains and chemical reaction networks. SIAM Journal on Applied Mathematics, 83(4):1444–1466, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M150770X.

Li:2023:FHB

[390] Sirui Li and Jie Xu. Frame hydrodynamics of biaxial nematics from molecular-theory-based tensor models. SIAM Journal on Applied Mathematics, 83(4):1467–1495, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/21M1465792.

Jin:2023:IPS

[391] Bangti Jin, Yavar Kian, and Zhi Zhou. Inverse problems for subdiffusion from observation at an unknown terminal time. SIAM Journal on Applied Mathematics, 83(4):1496–1517, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1529105.

Bressloff:2023:RES

[392] Paul C. Bressloff. Renewal equations for single-particle diffusion in multilayered media. SIAM Journal on Applied Mathematics, 83(4):1518–1545, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

                                        tronic). URL https://epubs.siam.
                                        org/doi/10.1137/23M1545835.

Li:2023:DPH

[393] Wei Li, Junshan Lin, and Hai Zhang. Dirac points for the honeycomb lattice with impenetrable obstacles. SIAM Journal on Applied Mathematics, 83(4):1546–1571, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1505116.

Jia:2023:ATD

[394] Chen Jia and Youming Li. Analytical time-dependent distributions for gene expression models with complex promoter switching mechanisms. SIAM Journal on Applied Mathematics, 83(4):1572–1602, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M147219X.

Lyu:2023:PSE

[395] Jhih-Hong Lyu and Tai-Chia Lin. PBsteric equations: a general model of Poisson–Boltzmann equations. SIAM Journal on Applied Mathematics, 83 (4):1603–1622, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://

                                        epubs.siam.org/doi/10.1137/22M1516270.

Ablowitz:2023:UAF

[396] Mark J. Ablowitz, Justin T. Cole, and S. D. Nixon. Unified approach to Floquet lattices, topological insulators, and their nonlinear dynamics. SIAM Journal on Applied Mathematics, 83(4):1623– 1653, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/22M1496979.

Li:2023:FMI

[397] Jiwei Li, Lingyun Qiu, Zhongjing Wang, and Hui Yu. Flow measurement: an inverse problem formulation. SIAM Journal on Applied Mathematics, 83(4):1654–1676, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1530720.

Kraisler:2023:TMW

[398] Joseph Kraisler, Wei Li, Kui Ren, John C. Schotland, and Yimin Zhong. Transport models for wave propagation in scattering media with nonlinear absorption. SIAM Journal on Applied Mathematics, 83(4):1677–1695, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1533505.

Jahedi:2023:SSN

[399] Sana Jahedi, Timothy Sauer, and James A. Yorke. Structured systems of nonlinear equations. SIAM Journal on Applied Mathematics, 83(4):1696–1716, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1529178.

Craciun:2023:AFW

[400] Gheorghe Craciun, Jiaxin Jin, and Polly Y. Yu. An algorithm for finding weakly reversible deficiency zero realizations of polynomial dynamical systems. SIAM Journal on Applied Mathematics, 83(4):1717–1737, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1499558.

Ghandriche:2023:SRO

[401] Ahcene Ghandriche and Mourad Sini. Simultaneous reconstruction of optical and acoustical properties in photoacoustic imaging using plasmonics. SIAM Journal on Applied Mathematics, 83(4):1738–1765, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL https://epubs.siam.org/doi/
                                        10.1137/22M1534730.

Skardal:2023:PFO

[402] Per Sebastian Skardal. Pattern formation and oscillations in nonlinear random walks on networks. SIAM Journal on Applied Mathematics, 83(5):1767– 1784, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

                                        siam.org/doi/10.1137/22M1536662.

Jiang:2023:PMS

[403] Xue Jiang, Maohui Lyu, Tao Yin, and Weiying Zheng. A PML method for signal-propagation problems in axon. SIAM Journal on Applied Mathematics, 83(5):1785–1805, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1544063.

Wang:2023:BRR

[404] Feng-Bin Wang, Lei Zhang, and Xiao-Qiang Zhao. Basic reproduction ratios for time-periodic homogeneous evolution systems. SIAM Journal on Applied Mathematics, 83(5):1806–1831, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1531865.

Ying:2023:MCS

[405] Fabian Ying, Alisdair O. G. Wallis, Mason A. Porter, Sam D. Howison, and Mariano Beguerisse-Díaz. Minimizing congestion in single-source, single-sink queueing networks. SIAM Journal on Applied Mathematics, 83(5):1832–1853, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/21M1457515.

Ren:2023:BLL

[406] Xiaofeng Ren and Juncheng Wei. The BCC lattice in a long-range interaction system. SIAM Journal on Applied Mathematics, 83(5):1854–1871, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1511722.

Cao:2023:UTP

[407] Fei Cao and Sébastien Motsch. Uncovering a two-phase dynamics from a dollar exchange model with bank and debt. SIAM Journal on Applied Mathematics, 83(5):1872–1891, ????

2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1518621.

Zhang:2023:BAT

[408] Yuyue Zhang, Liqi Xie, Yueping Dong, Jicai Huang, Shigui Ruan, and Yasuhiro Takeuchi. Bifurcation analysis in a tumor-immune system interaction model with dendritic cell therapy and immune response delay. SIAM Journal on Applied Mathematics, 83(5):1892– 1914, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/22M1533979.

Li:2023:RME

[409] Jialei Li, Xiaodong Liu, and Qingxiang Shi. Reconstruction of multiscale elastic sources from multifrequency sparse far field patterns. SIAM Journal on Applied Mathematics, 83(5):1915–1934, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1544257.

Dipierro:2023:ALF

[410] Serena Dipierro, Giovanni Giacomin, and Enrico Valdinoci. Analysis of the Lévy flight foraging hypothesis in R^{n} and unreliability of the most rewarding strategies. SIAM Journal on Applied Mathematics, 83(5):1935–1968, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1526563.

Parkinson:2023:ARD

[411] Christian Parkinson and Weinan Wang.

Analysis of a reaction-diffusion SIR epidemic model with noncompliant behavior. SIAM Journal on Applied Mathematics, 83(5):1969–2002, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/23M1556691.

Meng:2023:DDB

[412] Shixu Meng. Data-driven basis for reconstructing the contrast in inverse scattering: Picard criterion, regularity, regularization, and stability. SIAM Journal on Applied Mathematics, 83(5):2003– 2026, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL https://epubs.

siam.org/doi/10.1137/23M1545409.

Cebra:2023:SSC

[413] Christopher Cebra and Alexander Strang. Similarity suppresses cyclicity: Why similar competitors form hierarchies. SIAM Journal on Applied Mathematics, 83(5):2027–2051, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

URL https://epubs.siam.org/doi/
10.1137/22M1503099.

Piu:2023:DSA

[414] Matteo Piu, Michael Herty, and Gabriella Puppo. Derivation and stability analysis of a macroscopic multilane model for traffic flow. SIAM Journal on Applied Mathematics, 83(5):2052–2072, ???? 2023. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (elec-

tronic). URL https://epubs.siam.
org/doi/10.1137/22M1543288.